### v1

# On a fast consistent selection of nested models with possibly unnormalised probability densities

Abstract
Models with unnormalized probability density functions are ubiquitous in statistics, artificial intelligence and many other fields. However, they face significant challenges in model selection if the normalizing constants are intractable. Existing methods to address this issue often incur high computational costs, either due to numerical approximations of normalizing constants or evaluation of bias corrections in information criteria. In this paper, we propose a novel and fast selection criterion, T-GIC, for nested models, allowing direct data sampling from a possibly unnormalized probability density function. T-GIC gives a consistent selection under mild regularity conditions and is computationally efficient, benefiting from a multiplying factor that depends only on the sample size and the model complexity. Extensive simulation studies and real-data applications demonstrate the efficacy of T-GIC in the selection of nested models with unnormalized probability densities.
Keywords: unnormalized probability densities, gradient-based information criterion, consistent model selection, computational efficiency, nested models
1 Introduction
Models with unnormalized probability density functions are ubiquitous. In statistics, artificial intelligence, statistical mechanics and many other fields, often we only want or are able to stipulate the general shape of the models’ distributions without requiring their underlying probability density functions (PDFs) to integrate to unity, hence the notion of unnormalized PDFs. In such cases, the normalizing constants are either extremely difficult or impossible to compute directly. They arise in various circumstances, e.g. non-conjugacy in Bayesian posteriors (Andrade and Rathie, 2017), partition functions in statistical mechanics (Frigg and Werndl, 2024), directional distributions of data on a sphere (Pewsey and García-Portugués, 2021), Ising model in spatial statistics (Friel, 2013), and distributions with high-dimensional latent variables (Murray and Salakhutdinov, 2008). In short, unnormalized PDFs pose a significant challenge in likelihood-based model comparison. Addressing the challenge, we focus on the selection of nested models with possibly unnormalized PDFs.
Model selection involves choosing the best statistical model from several candidates based on the observed data, with each candidate having possibly a different number of parameters (Rao et al., 2001). Among the forerunners and the most prominent members of this vast industry, Akaike (1974) proposed the Akaike Information Criterion (AIC) based on the Kullback-Leibler divergence that is underpinned by the Shannon entropy. Here, the model with the minimum AIC is selected as the best model. As the other, Schwarz (1978) developed the Bayesian Information Criterion (BIC) via a Laplace approximation of the Bayes factor. Here, the model with the minimum BIC value is selected. Unfortunately, none of them is applicable for the selection of models with unnormalized PDFs.
At present, there are essentially two main approaches to overcome the challenge. One approach is to first estimate the intractable normalizing constant by numerical approximations or Markov Chain Monte Carlo. See, for example, (Baker, 2022) and (MacKay, 2003; Congdon, 2006). While simple and intuitive, this approach is computationally intensive, especially for random vectors in high-dimensional settings. Another approach is to avoid the calculation of the normalizing constant altogether. For this, the score matching method (Hyvärinen, 2005) provides significant potential through efficient computation, as it does not depend on the normalizing constant. This method has been applied to Bayesian model selection (Dawid and Musio, 2015; Shao et al., 2019) that focuses mainly on the issue of improper priors while the data are still sampled from a normalized distribution. In this approach, consistent selection is only proved for non-nested models ((Shao et al., 2019), p.1826), with the situation for nested models, such as polynomial regression models or autoregressive (AR) models, remaining unclear. Our approach is different. We allow the data to be sampled directly from an unnormalized PDF.
Recently, Matsuda et al. (2021) proposed an information criterion for selection of models with unnormalized PDFs estimated via noise contrastive estimation (NCE) or score matching. Slightly later, Cheng and Tong (2024) introduced a new entropy, the G-entropy, that includes an information criterion for model selection via a negative Hyvärinen score. In the above information criteria, bias correction is involved and the estimation of the bias term can be computationally intensive. For example, with parameter dimension $k$ and sample size $n$, matrix calculations for the bias in Cheng and Tong (2024) run in $O(nk^{2}+k^{3})$ time. Similar order applies to Matsuda et al. (2021)
In this paper, we propose a fast information criterion for consistent selection of nested models with unnormalized PDFs. We name the criterion T-GIC. In it, we introduce a multiplying factor $C(n,k)$, which depends only on the sample size $n$ and the order $k$ of the candidate model $M_{k}$, resulting in a drastic reduction of computation to $O(1)$. We show that T-GIC achieves consistent model selection for strictly nested models under mild regularity conditions. Further, we demonstrate the efficacy of T-GIC through simulation on AR and polynomial regression models with unnormalized PDFs as well as a model with a bivariate von Mises PDF with bounded support. Finally, we apply the T-GIC to real data from diverse domains, including finance, automotive engineering, and wind direction analysis.
This paper is organized as follows. In Section 2, we provide a brief review of F-divergence and associated notions. In Section 3, we introduce the T-GIC and prove its consistency for model selection under regularity assumptions. In Section 4, we conduct simulations to demonstrate the efficacy of T-GIC. In Section 5, we apply T-GIC to various real-world data, offering some insights. In Section 6, we conclude with a discussion of our findings and potential directions for future research. Proofs are provided in Appendix A.
2 A New Information Criterion
It is well known that the Kullback-Leibler divergence is underpinned by the Shannon entropy. Recently, Cheng and Tong (2024) have developed a new entropy that parallels Shannon’s. It is called the G-entropy (sometimes shortened to Gentropy). They have shown how it underpins the Fisher divergence and leads to the GIC for model selection.
2.1 Fisher Divergence and Gentropy
Let $p(x)$ be a PDF on $\mathbb{R}^{d}$ under the following assumptions:
Assumption 1.
$p(x)$ is twice differentiable on $\mathbb{R}^{d}$;
Assumption 2.
$p(x)$, $\nabla_{x}p(x)$ and $\nabla_{x}\nabla_{x}p(x)$ are all square-integrable on $\mathbb{R}^{d}$;
Assumption 3.
For every $x\in R^{d}$ with $x=(x_{1},\cdots,x_{d})$ and for each boundary point of $x_{i},~{}i=1,\cdots,d$,
$p(x_{1},\cdots,x_{i-1},-\infty,x_{i+1},\cdots,x_{d})\equiv 0\text{ and }p(x_{1},\cdots,x_{i-1},+\infty,x_{i+1},\cdots,x_{d})\equiv 0.$
Let $q(x)$ be a PDF on $\mathbb{R}^{d}$ satisfying Assumptions 1 and 2. The Fisher divergence between PDF $p$ and PDF $q$ is defined as
$D_{F}(p||q)=E_{p}||\nabla_{x}\log p(x)~{}-~{}\nabla_{x}\log q(x)||^{2}.$ (2.1)
The objective function $W$ of PDF $q$ at point $x$ is a negative Hyvärinen score (Hyvärinen, 2005), defined by
$W(x,q)=-||\nabla_{x}\log q(x)||^{2}~{}-~{}2\Delta_{x}\log q(x),$ (2.2)
where $\Delta$ denotes the Laplacian, i.e. $\Delta_{x}f(x)=\sum_{i=1}^{d}\frac{\partial^{2}f(x)}{\partial^{2}x_{i}}$.
Next, within their new entropy structure, Cheng and Tong (2024) define, among other notions, Gentropy $H_{G}(p)$ and G-cross-entropy $H_{G}(p,q)$ as
$H_{G}(p)=E_{p}||\nabla_{x}\log p(x)||^{2},$ (2.3)
$H_{G}(p,q)=E_{p}[W(x,q)].$ (2.4)
Integration by parts yields the following relationship
$D_{F}(p||q)=H_{G}(p)-H_{G}(p,q).$
Denote $x$ and $y$ as two random vectors with joint PDF $p(x,y)$. Let $\theta$ be the unknown $h$-dimensional parameter vector for $p(x)$, $h\geq 1$. Tables 1 and 2 summarize the comparison between Shannon entropy and G-entropy across different definitions and properties.
Table 1: Comparison between Shannon entropy and G-entropy: definitions.
Definitions Shannon entropy G-entropy
Objective function $Loss(p(x))=-\log p(x)$ $W(x,p)=-||\nabla_{x}\log p(x)||^{2}~{}-~{}2\Delta_{x}\log p(x)$
Entropy $H(x)=H(p)=E_{p}[Loss(p(x))]$ $H_{G}(x)=H_{G}(p)=E_{p}||\nabla_{x}\log p(x)||^{2}$
Cross entropy $H(p,q)=E_{p}[Loss(q(x))]$ $H_{G}(p,q)=E_{p}[W(x,q)]$
Joint entropy $H(x,y)=H(p_{(x,y)})$ $H_{G}(x,y)=H_{G}(p_{(x,y)})$
Conditional entropy $H(y|x)=E_{p_{(x,y)}}[Loss(\frac{p(x,y)}{p(x)})]$ $H_{G}(y|x)=E_{p(x,y)}[W((x,y),\frac{p(x,y)}{p(x)})]$
Mutual information $I(x,y)=H(x)+H(y)-H(x,y)$ $I_{G}(x,y)=H_{G}(x,y)-H_{G}(x)-H_{G}(y)$
Information matrix $IM_{\theta}=E_{p}[[\nabla_{\theta}\log p(x)][\nabla_{\theta}\log p(x)]^{\tau}]$ $GIM_{x}=E_{p}[[\nabla_{x}\log p(x)][\nabla_{x}\log p(x)]^{\tau}]$
Divergence $D_{KL}(p||q)=E_{p}[Loss(\frac{q(x)}{p(x)})]$ $=H(p,q)-H(p)$ $D_{F}(p||q)=E_{p}||\nabla_{x}\log p(x)~{}-~{}\nabla_{x}\log q(x)||^{2}$ $=H_{G}(p)-H_{G}(p,q)$
Table 2: Comparison between Shannon entropy and G-entropy: properties.
Shannon entropy G-entropy
1 $H(p,p)=H(p)=E_{p}[Loss(p(x))]$ $H_{G}(p,p)=H_{G}(p)=E_{p}[W(x,p)]$
2 $H(p)\leq H(p,q)$ $H_{G}(p)\geq H_{G}(p,q)$
3 $\min_{q}\{H(p,q)\}=H(p)$ $\max_{q}\{H_{G}(p,q)\}=H_{G}(p)$
4 $H(x,y)\leq H(x)+H(y)$, with equality iff $x$ and $y$ are independent. $H_{G}(x,y)\geq H_{G}(x)+H_{G}(y)$, with equality iff $x$ and $y$ are independent.
5 $H(y|x)=H(x,y)-H(x)$ $H_{G}(y|x)=H_{G}(x,y)-H_{G}(x)$
2.2 GIC Estimate
Following equation (2.1), given PDF $p$, minimizing $D_{F}(p||q)$ with respect to $q$ is equivalent to maximizing the G-cross-entropy $H_{G}(p,q)$. When $p$ is the true data PDF and $q$ the model PDF, $H_{G}(p,q)$ inspires the maximum GIC estimate (MGICE).
Specifically, let $p_{x}$ be the data PDF, and $p_{M(\theta)}$ be a PDF of model $M$ with unknown $h$-dimensional parameter vector $\theta$ ($h\geq 1$) in parameter space $\Theta$. The true parameter $\theta^{*}$ can be obtained by
$\theta^{*}=\arg\min_{\theta\in\Theta}D_{F}(p_{x}||p_{M(\theta)})=\arg\max_{\theta\in\Theta}H_{G}(p_{x},p_{M(\theta)}).$ (2.5)
Let $x_{1},\cdots,x_{n}$ be a sample in $\mathbb{R}^{d}$ from the data PDF $p_{x}$. An unbiased estimate of $H_{G}(p_{x},p_{M(\theta)})$ is given by
$GIC(M(\theta))=\frac{1}{n}\sum_{i=1}^{n}W(x_{i},p_{M(\theta)}).$ (2.6)
Hence, the MGICE of $\theta$ is as follows
$\hat{\theta}=\hat{\theta}_{n}=arg\max_{\theta\in\Theta}\{GIC(M(\theta))\}.$ (2.7)
2.3 GIC for Model Selection
Consider a collection of candidate parametric models $M_{1},\cdots,M_{K}$, denoted as $M_{k}(\theta_{k}),k=1,\cdots,K$, with $\theta_{k}$ ($h_{k}\geq 1$) being an unknown $h_{k}$-dimensional parameter vector in parameter space $\Theta.$ Similar to AIC, an unbiased $GIC_{c}$ for model selection (Cheng and Tong 2024) is derived under mild regularity conditions by correcting the bias, $B$, in $nGIC$ as follows:
$GIC_{c}(M_{k}(\hat{\theta}_{k}))=GIC_{n}(M_{k}(\hat{\theta}_{k}))-B_{k},$ (2.8)
where $\hat{\theta}_{k}$ is the MGICE of $\theta_{k}$, $GIC_{n}(M_{k}(\hat{\theta}_{k}))=nGIC(M_{k}(\hat{\theta}_{k}))$, and
$B_{k}$ $=-tr\{E_{p_{x}}[\nabla_{\theta}W(x,p_{M_{i}(\theta^{*})})\nabla_{\theta}^{T}W(x,p_{M_{i}(\theta^{*})})]E_{p_{x}}^{-1}[\nabla_{\theta}^{2}W(x,p_{M_{i}(\theta^{*})})]\}.$ (2.9)
By maximizing $GIC_{c}$, an appropriate model from $M_{1},\cdots,M_{K}$ is selected.
3 A New Model Selection Criterion: T-GIC
In this section, we propose a fast model selection criterion, T-GIC, and show the consistency for a finite sequence of strictly nested models under mild regularity conditions.
3.1 A Fast Model Selection
Let $x_{1},\cdots,x_{n}$ be a sample in $\mathbb{R}^{d}$ from the data PDF $p_{x}$. Suppose we have a collection of candidate parametric models as described in Section 2.3. The $T\text{-}GIC(k)$ of model $M_{k}$ is defined as
$T\text{-}GIC(k)=C(n,k)GIC(M_{k}(\hat{\theta}_{k})),$ (3.1)
where $C(n,k)$ is a constant depending only on $n$ and $k$, $GIC(M_{k}(\hat{\theta}_{k}))$ is defined by equation (2.6), and $\hat{\theta}_{k}$ is the MGICE of $\theta_{k}$ for model $M_{k}$.
We propose to select the model that maximizes $T\text{-}GIC(k)$. Compared to $GIC_{c}$ in equation (2.8), $T\text{-}GIC$ significantly reduces the computational costs by introducing a factor $C(n,k)$ to bypass the bias correction calculation. Note that estimating the bias, B, based on a sample of size $n$ involves a calculation $O(nh_{k}^{2}+h_{k}^{3})$, which becomes computationally expensive when $h_{k}$ or $n$ is large. In contrast, using the multiplying factor $C(n,k)$ reduces the computational cost to only $O(1)$. The following section shows the consistency of the proposed model selection criterion T-GIC.
3.2 Consistency of T-GIC
We first introduce several assumptions, which have been used by Song et al. (2020) and Cheng and Tong (2024).
Assumption 4.
$p_{x}=p_{M(\theta^{*})},$ where $\theta^{*}$ is the true parameter in $\Theta$. Furthermore, $p_{M(\theta)}\neq p_{M(\theta^{*})}$ whenever $\theta\neq\theta^{*}$.
Assumption 5.
$p_{M(\theta)}(x)>0,~{}\forall\theta\in\Theta$ and $\forall x$.
Assumption 6.
The parameter space $\Theta$ is compact.
Assumption 7.
Both $\nabla_{x}^{2}\log p_{M(\theta)}(x)$ and $[\nabla_{x}\log p_{M(\theta)}(x)][\nabla_{x}\log p_{M(\theta)}(x)]^{T}$ are Lipschitz continuous in respect of Frobenious norm. Specifically, $\forall\theta_{1},\theta_{2}\in\Theta$,
$||\nabla_{x}^{2}\log p_{M(\theta_{1})}(x)-\nabla_{x}^{2}\log p_{M(\theta_{2})}(x)||_{F}\leq L_{1}(x)||\theta_{1}-\theta_{2}||_{2},$
and
$||[\nabla_{x}\log p_{M(\theta_{1})}(x)][\nabla_{x}\log p_{M(\theta_{1})}(x)]^{T}-[\nabla_{x}\log p_{M(\theta_{2})}(x)][\nabla_{x}\log p_{M(\theta_{2})}(x)]^{T}||$
$\leq L_{2}(x)||\theta_{1}-\theta_{2}||_{2}.$
In addition, $E_{p_{x}}[L_{1}^{2}(x)]<\infty$ and $E_{p_{x}}[L_{2}^{2}(x)]<\infty.$
Assumption 8.
For $\theta_{1}$, $\theta_{2}$ near $\theta^{*}$, and $\forall i,j$,
$||\nabla_{\theta}^{2}\partial_{i}\partial_{j}\log p_{M(\theta_{1})}-\nabla_{\theta}^{2}\partial_{i}\partial_{j}\log p_{M(\theta_{2})}||_{F}\leq M_{i,j}(x)||\theta_{1}-\theta_{2}||_{2},$
and
$||\nabla_{\theta}^{2}\partial_{i}\log p_{M(\theta_{1})}\partial_{j}\log p_{M(\theta_{1})}-\nabla_{\theta}^{2}\partial_{i}\log p_{M(\theta_{1})}\partial_{j}\log p_{M(\theta_{2})}||_{F}\leq N_{i,j}(x)||\theta_{1}-\theta_{2}||_{2}.$
Here, $\partial_{i}$ refers to the partial derivative with respect to the component $x_{i}$ in the random vector $x=(x_{1},\cdots,x_{d}).$
Note that Assumptions 4 and 6 are standard conditions for proving the consistency of the maximum likelihood estimation (MLE). Assumption 5 is also used by Hyvärinen (2005). Assumption 7 defines Lipschitz continuity, while Assumption 8 describes Lipschitz smoothness for second derivatives. Based on these assumptions, Proposition 6 in Cheng and Tong (2024) shows the asymptotic normality of the MGICE, as stated in Lemma 1.
Lemma 1.
Under Assumptions 1-8 and let $\hat{\theta}_{n}$ be the MGICE, we have
$\sqrt{n}(\hat{\theta}_{n}-\theta^{*})\xrightarrow{dist}N(0,D^{-1}(\theta^{*})\Lambda(\theta^{*})D^{-1}(\theta^{*})),$ (3.2)
where
$D(\theta^{*})=-E_{p_{x}}[\nabla_{\theta}^{2}W(x,p_{M(\theta^{*})})],$ (3.3)
$\Lambda(\theta^{*})=E_{p_{x}}[\nabla_{\theta}W(x,p_{M(\theta^{*})})\nabla_{\theta}^{T}W(x,p_{M(\theta^{*})})].$ (3.4)
Moreover, Table 3 provides a detailed comparison between MLE and MGICE.
Table 3: Comparison between MLE and MGICE.
Aspects MLE MGICE
Objective function $\log[p_{M(\theta)}(x_{i})]$ $W(x_{i},p_{M(\theta)})=-||\nabla_{x}\log p_{M(\theta)}(x_{i})||^{2}~{}-~{}2\Delta_{x}\log p_{M(\theta)}(x_{i})$
Sample function $l_{n}(M(\theta))=\sum_{i=1}^{n}\log[p_{M(\theta)}(x_{i})]$ $GIC_{n}(M(\theta))=\sum_{i=1}^{n}W(x_{i},p_{M(\theta)})$
Estimate $\hat{\theta}_{MLE}=arg\max_{\theta\in\Theta}\{l_{n}(M(\theta))\}$ $\hat{\theta}_{GIC}=arg\max_{\theta\in\Theta}\{GIC_{n}(M(\theta))\}$
Information 1 $I(\theta^{*})=E_{p_{x}}[\nabla_{\theta}\log[p_{M(\theta^{*})}(x)]\nabla_{\theta}^{T}\log[p_{M(\theta^{*})}(x)]]$ $\Lambda(\theta^{*})=E_{p_{x}}[\nabla_{\theta}W(x,p_{M(\theta^{*})})\nabla_{\theta}^{T}W(x,p_{M(\theta^{*})})]$
Information 2 $J(\theta^{*})=-E_{p_{x}}[\nabla_{\theta}^{2}\log[p_{M(\theta^{*})}(x)]]$ $D(\theta^{*})=-E_{p_{x}}[\nabla_{\theta}^{2}W(x,p_{M(\theta^{*})})]$
Consistency $\hat{\theta}_{MLE}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\theta^{*}~{}\text{ as }n\longrightarrow\infty$ $\hat{\theta}_{GIC}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\theta^{*}~{}\text{ as }n\longrightarrow\infty$
Asymptotics $\sqrt{n}(\hat{\theta}_{MLE}-\theta^{*})\xrightarrow{dist}N(0,J^{-1}(\theta^{*})I(\theta^{*})J^{-1}(\theta^{*}))$ $\sqrt{n}(\hat{\theta}_{GIC}-\theta^{*})\xrightarrow{dist}N(0,D^{-1}(\theta^{*})\Lambda(\theta^{*})D^{-1}(\theta^{*}))$
Next, we claim that under suitable regularity conditions, $n\times[\log GIC(k)-\log GIC(k_{0})]$ converges weakly to some non-negative distribution, where $k_{0}$ is the smallest $k$ such that $M_{k}$ contains the true model.
Proposition 1.
Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, i.e., $M_{k}\subsetneq M_{k+1}$ for all $k=1,\cdots,K-1$, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model. Consider the model $M=M_{k}$ for some $k>k_{0}$ with its parameter $\theta$ partitioned into two sub-vectors $\alpha$ and $\beta$ such that the true parameter $\theta^{*}$ obtains when $\alpha=\alpha^{*}$, $\beta=\beta^{*}=0$. Similarly, partition $D(\theta^{*})$ defined in (3.3) into a 2 by 2 block matrix:
$D(\theta^{*})=\begin{pmatrix}D(\alpha^{*},\alpha^{*})&D(\alpha^{*},\beta^{*})\\ D(\beta^{*},\alpha^{*})&D(\beta^{*},\beta^{*})\end{pmatrix}.$ (3.5)
Assume that $M_{k_{0}}$ is obtained from $M$ by constraining $\beta=0$. Under Assumptions 1-8, then
$n\times[\log GIC(k)-\log GIC(k_{0})]\xrightarrow{dist}Z^{T}A(\theta^{*})Z,$ (3.6)
and
$A(\theta^{*})=-4H_{G}^{2}(p^{*})B^{T}(\theta^{*})\{D(\beta^{*},\beta^{*})-D(\beta^{*},\alpha^{*})D^{-1}(\alpha^{*},\alpha^{*})D(\alpha^{*},\beta^{*})\}B(\theta^{*}),$ (3.7)
$B(\theta^{*})=\{D(\beta^{*},\beta^{*})-D(\beta^{*},\alpha^{*})D^{-1}(\alpha^{*},\alpha^{*})D(\alpha^{*},\beta^{*})\}^{-1}$
$\begin{pmatrix}-D(\beta^{*},\alpha^{*})D^{-1}(\alpha^{*},\alpha^{*})&I\end{pmatrix}\Lambda^{1/2}(\theta^{*}),$ (3.8)
where $Z$ is a $(k-k_{0})$-dimensional standard normal random vector, $p^{*}=p_{M(\alpha^{*},\beta^{*})}$, and $I$ denotes the identity matrix of dimension $k_{0}$.
Now, we establish the consistency of our T-GIC method.
Theorem 1.
Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model. Suppose the following conditions are satisfied:
  1. (1)
    For any $K\geq k\geq k_{0}$, Assumptions 4-8 hold and the MGICE of their parameter is $\sqrt{n}$-consistent with an asymptotic normal distribution whose covariance matrix is invertible.
  2. (2)
    For any $1\leq k<k_{0}$, Assumptions 6-7 hold.
  3. (3)
    $C(n,k)$ is such that (i) for any $k$, $C(n,k)\to 1$ as $n\to\infty$ and (ii) for any $K\geq k_{1}>k_{2}\geq 1$, $n\times\log\{C(n,k_{1})/C(n,k_{2})\}\to-\infty$ as $n\to\infty$.
Let $\hat{k}=\arg\max_{1\leq k\leq K}T\text{-}GIC(k)$. Then $\hat{k}$ converges to $k_{0}$, in probability.
Remark 1.
For ease of exposition, we have, so far, assumed that the data are independent and identically distributed (IID). However, Proposition 1 and Theorem 1 can be extended to dependent data, under suitable regularity conditions. For instance, for stationary ergodic finite-order homogeneous Markov processes including autoregressive models, we can generalize GIC as follows, for an order-$k$ Markov process:
$CGIC(M(\theta))=\frac{1}{n-k}\sum_{i=k+1}^{n}W(x_{i},p_{M(\theta)}(\cdot|x_{i-1},\cdots,x_{i-k})),$ (3.9)
where $p_{M(\theta)}(\cdot|x_{i-1},\cdots,x_{i-k})$ is the conditional pdf of $x_{i}$ given its lags 1 to $k$. It follows from Assumptions 1–3 that $\nabla_{\theta}CGIC(M(\theta))$, evaluated at the true parameter, is a martingale difference sequence. By appealing to some variant of the martingale central limit theorem (Hall and Heyde, 2014, Chapter 3), the proofs of Proposition 1 and Theorem 1 for the IID case can be easily extended to the case of stationary ergodic finite-order Markov processes, with MGICE and T-GIC modified accordingly.
Mimicking AIC and BIC, we set $C(n,k)$ to $\exp\{-2\#(M_{k})/n\}$ and $n^{-\#(M_{k})/n}$, respectively, where #$\#(M_{k})$ is the number of independently adjusted parameters in model $M_{k}$. From Theorem 1, the following corollaries follow immediately.
Corollary 1.
Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model. If $K=k_{0}$, under Assumptions 1-8, then $T\textbf{-}GIC1=\exp\{-2\#(M_{k})/n\}GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
Corollary 2.
Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model. Under Assumptions 1-8, then $T\textbf{-}GIC2=n^{-\#(M_{k})/n}GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
However, similar to AIC, T-GIC1 may fail to provide consistent selection when $K>k_{0}$. In the following sections, we will compare the performance of T-GIC1 and T-GIC2 through simulations and applications.
4 Simulation Study
In this section, we assess the efficacy of T-GIC1 and T-GIC2 for models with unnormalized PDFs. In Section 4.1, we study the consistency of MGICE for $N\times t$ PDF. In Sections 4.2 and 4.3, we evaluate the consistency (or otherwise) of T-GIC1 and T-GIC2 and the MGICE for two true models, namely the AR model and the polynomial regression model, each defined on $\mathbb{R}$ and driven by $N\times t$ noise/errors. In Section 4.4, we examine a model with a bvariate von Mises PDF on bounded support. Upon replications, the empirical distribution of selected orders by T-GIC’s are obtained, as well as the average MGICE and its standard deviation (SD) for parameters in the true model.
4.1 $N\times t$ PDF
First, we evaluate the efficacy of the MGICE for unnormalized PDFs. Consider an $N\times t(\alpha,k)$ PDF defined as the product of a normal PDF and a Student t-PDF, given by
$f_{N\times t}{(x\mid\alpha,k)}=C(\theta)\frac{\exp\{-\frac{\alpha x^{2}}{2}\}}{(1+x^{2})^{k}},$ (4.1)
for $-\infty<x<\infty$, where $\alpha>0$, $k>0$, $\theta=(\alpha,k)^{T}$ and $C(\theta)$ is the normalizing constant. $N\times t$ is fat-tailed, with the parameters $\alpha$ and $k$ collectively controlling the tail behavior. Specifically, $\alpha$ represents the scale parameter, and $k$ is the power parameter, which is related to the degrees of freedom $\nu$ of a t-PDF via $k=(\nu+1)/2$.
The normalizing constant $C(\theta)$ is generally computationally intractable, except in certain special cases, such as when $k$ is an integer. Prior to the MLE, Baker (2022) either (i) derived $C(\theta)$ under the above condition, or (ii) used numerical quadrature to approximate $C(\theta)$ for general $k$. Instead of the MLE, we employ MGICE for data fitting, which circumvents the need for normalizing constants.
Considering the affine transformation $X=(Y-\mu)/s$ in Baker (2022), the PDF of $Y$ is $\frac{1}{s}f_{N\times t}(\frac{y-\mu}{s}|\alpha,k)$, with 4 parameters $(\mu,s,\alpha,k)$, where $-\infty<\mu<\infty$ and $s>0$. Let $\{y_{i},i=1,2,\cdots,n\}$ represent $n$ observations and denote $f_{y_{i}}(\theta)=\frac{1}{s}f_{N\times t}(\frac{y_{i}-\mu}{s}|\alpha,k)$. Following routine derivations, we have
$\nabla_{y_{i}}\log(f_{y_{i}})$ $=-\frac{1}{s}[\alpha x_{i}+\frac{2kx_{i}}{1+x_{i}^{2}}],$ (4.2)
and
$\Delta_{y_{i}}\log(f_{y_{i}})$ $=-\frac{1}{s^{2}}[\alpha+\frac{2k(1-x_{i}^{2})}{(1+x_{i}^{2})^{2}}],$ (4.3)
where $x_{i}=(y_{i}-\mu)/s$. From equations (2.2), (2.6) and (2.7), we obtain the MGICE, $\hat{\theta}_{GIC}$.
We conduct 100 replications to evaluate the performance of the MGICE, with sample size $n=1000,3000,\text{and }5000$, where the true parameter values are set to $(\mu^{*},s^{*},\alpha^{*},k^{*})=(0.3,0.5,0.5,1.5)$. Samples are generated using the acceptance-rejection method, by generating random numbers from the $N(0,1/\alpha)$ distribution. For numerical optimization in MGICE, we use the Adaptive Moment Estimation (Adam) algorithm to jointly optimize all parameters. Similar to Baker (2022), for all experiments involving the $N\times t$ PDF, we use the sample mean and standard deviation as starting values for $\mu$ and $s$, and regular initial values for $\alpha$ and $k$, e.g. $\alpha=0.25$ or 1, and $k=1$ or 2. Table 4 presents the average MGICE and its SD, showing good overall consistency of MGICE, although there is still room for improvement for the parameters $s^{*},\alpha^{*},$ and $k^{*}$.
Table 4: The average MGICE and its SD for $N\times t$ PDF.
$N$ Parameter MGICE (SD)
$\mu$ $s$ $\alpha$ $k$
1000 0.30 (0.02) 0.52 (0.10) 0.52 (0.25) 1.62 (0.39)
3000 0.30 (0.01) 0.53 (0.06) 0.48 (0.14) 1.68 (0.19)
5000 0.30 (0.01) 0.53 (0.04) 0.48 (0.10) 1.70 (0.14)
4.2 AR Model with $N\times t$ Noise
Now, we consider an AR model with $N\times t$ distributed noise and use T-GIC method to select the model order. The stationary mean-centered AR model of order p with $N\times t$ noise is given by
$X_{t}-c=a_{1}(X_{t-1}-c)+\cdots+a_{p}(X_{t-p}-c)+s\varepsilon_{t},$ (4.4)
where $-\infty<a_{1},\cdots,a_{p},c<\infty$, $s>0$, and $\varepsilon_{t}\sim N\times t(\alpha,k),$ identically and independently.
Let $\{x_{t},t=1,2,\cdots,N\}$ denote $N$ observations from the above model and denote the parameter as $\theta=(a_{1},\cdots,a_{p},c,s,\alpha,k)^{T}$. Let $y_{t}=x_{t}-c$ and $\mu_{t}=a_{1}y_{t-1}+\cdots+a_{p}y_{t-p}$. Since $\varepsilon_{t}=(y_{t}-\mu_{t})/s$, the conditional density of $X_{t}$ is $f_{x_{t}}=f_{p}(x_{t}\mid x_{t-1},\cdots,x_{t-p},\theta)=\frac{1}{s}f_{N\times t}(\frac{y_{t}-\mu_{t}}{s}|\alpha,k)$. From equations (4.2) and (4.3), we have $\nabla_{x_{t}}\log(f_{x_{t}})$ and $\Delta_{x_{t}}\log(f_{x_{t}})$.
Suppose we have a collection of candidate AR(p) models for $p=1,\cdots,L$, $L$ being the maximum possible order. Then, $GIC_{N}$ is given by
$GIC_{N}(\theta)=\sum_{t=1}^{L}W(x_{t},f_{x_{t}}(\theta))+\sum_{t=L+1}^{N}W(x_{t},f_{x_{t}}(\theta)).$ (4.5)
where $W(x_{t},f_{x_{t}})=-||\nabla_{x_{t}}\log(f_{x_{t}})||^{2}-2\Delta_{x_{t}}\log(f_{x_{t}})$.
Discarding the first sum because $W(x_{t},f_{x_{t}}(\theta))$ for $t=1,\cdots L$ are unavailable in the AR(L) model and denoting $n=N-L$, we have
$GIC(\theta)$ $=\frac{1}{n}\sum_{t=L+1}^{N}W(x_{t},f_{x_{t}}(\theta)).$ (4.6)
We conduct 100 replications to obtain the frequency distribution of the selected model orders, ranging from 1 to 10, using T-GIC1 and T-GIC2, with sample size $N=1000,3000,$ and $5000$. The true model order $p^{*}$ is set to 3, with true parameters $(a_{1}^{*},a_{2}^{*},a_{3}^{*},c,s,\alpha,k)$ specified at $(0.50,-0.25,0.10,3.00,0.50,0.50,1.50)$. In each replication, the first 200 data are discarded to ensure stationarity.
Table 5 shows that T-GIC2 tends to underestimate the model order for smaller $N$ but increasingly selects the correct order as $N$ grows. The underfitting for small $N$ is not unexpected due to the small value of $a_{3}^{*}$. In contrast, T-GIC1 overestimates the model order with high probability even for large $N$. Furthermore, Table 6 reports the average MGICE and its SD for the parameters in the true model, suggesting strong consistency, especially for the AR coefficients and the mean-centering parameter $c$, when $N$ is large.
Table 5: The frequency distribution of selected orders for the AR(p) model.
$N$ Method Selected model order $p$
1 2 3 4 5 6 7 8 9 10
1000 T-GIC1 0 3 22 13 11 1 6 13 14 17
T-GIC2 1 12 52 9 7 0 4 2 6 7
3000 T-GIC1 0 0 28 6 7 9 9 5 16 20
T-GIC2 0 1 81 11 0 4 1 0 0 2
5000 T-GIC1 0 0 29 10 8 7 9 10 7 20
T-GIC2 0 0 88 5 2 1 2 1 0 1
Table 6: The average MGICE and its SD for the AR(3) model with $N\times t$ noise.
$N$ Parameter MGICE (SD)
$a_{1}$ $a_{2}$ $a_{3}$ $c$ $s$ $\alpha$ $k$
1000 0.51 (0.13) -0.24 (0.11) 0.11 (0.13) 3.01 (0.11) 0.51 (0.14) 0.43 (0.26) 1.56 (0.61)
3000 0.50 (0.03) -0.26 (0.03) 0.10 (0.02) 3.00 (0.01) 0.51 (0.10) 0.45 (0.14) 1.61 (0.42)
5000 0.50 (0.02) -0.25 (0.02) 0.10 (0.02) 3.00 (0.01) 0.52 (0.07) 0.46 (0.10) 1.67 (0.31)
4.3 Polynomial Regression Model with $N\times t$ Distributed Errors
Next, we consider a polynomial regression model with $N\times t$ distributed errors and use T-GIC to select the model. The polynomial regression model of degree p is given by
$y=\beta_{1}x+\cdots+\beta_{p}x^{p}+c+s\varepsilon,$ (4.7)
where $-\infty<\beta_{1},\cdots,\beta_{p},c<\infty$, $s>0$, and $\varepsilon\sim N\times t(\alpha,k),$ identically and independently.
Let $\{(x_{i},y_{i}),i=1,2,\cdots,n\}$ denote $n$ observations from the above model and denote the parameter as $\theta=(\beta_{1},\cdots,\beta_{p},c,s,\alpha,k)^{T}$. Let $\mu_{i}=\beta_{1}x_{i}+\cdots+\beta_{p}x_{i}^{p}+c$ and $\varepsilon_{i}=(y_{i}-\mu_{i})/s$, The conditional density of $y_{i}$ is given by $f_{y_{i}}=f_{p}(y_{i})=f_{p}(y_{i}\mid x_{i},\theta)=\frac{1}{s}f_{N\times t}(\frac{y_{i}-\mu_{i}}{s}|\alpha,k).$ From equations (4.2) and (4.3), we have $\nabla_{y_{i}}\log(f_{y_{i}})$ and $\Delta_{y_{i}}\log(f_{y_{i}})$.
Suppose we have a collection of candidate models of degree $p$, for $p=1,\cdots,L$, $L$ being the maximum possible degree. For each model, from equations (2.2), (2.6), and (2.7), $GIC$ and its MGICE are obtained.
We conduct 100 replications to obtain the frequency distribution of the selected model degree $p$, ranging from 1 to 10, using T-GIC1 and T-GIC2, with sample size $n=1000,3000,$ and $5000$. The true model order $p^{*}$ is set to 3, with true parameters $(\beta_{1}^{*},\beta_{2}^{*},\beta_{3}^{*},c,s,\alpha,k)$ specified as $(-1.5,2.0,5.0,3.0,0.5,0.5,1.5)$.
Table 7 suggests that both T-GIC1 and T-GIC2 consistently select the correct degree in probability as $n$ increases. Moreover, Table 8 reports the average MGICE and its SD for the parameters in the true model, showing good consistency as $n$ increases, although at a slightly slower rate for $\alpha$.
Table 7: The frequency distribution of selected degrees for the polynomial regression model.
$n$ Method Selected model degree $p$
1 2 3 4 5 6 7 8 9 10
1000 T-GIC1 0 0 94 6 0 0 0 0 0 0
T-GIC2 0 0 97 3 0 0 0 0 0 0
3000 T-GIC1 0 0 96 4 0 0 0 0 0 0
T-GIC2 0 0 96 4 0 0 0 0 0 0
5000 T-GIC1 0 0 97 3 0 0 0 0 0 0
T-GIC2 0 0 97 3 0 0 0 0 0 0
Table 8: The average MGICE and its SD for the cubic polynomial regression model.
$n$ Parameter MGICE (SD)
$\beta_{1}$ $\beta_{2}$ $\beta_{3}$ $c$ $s$ $\alpha$ $k$
1000 -1.50 (0.03) 2.00 (0.01) 5.00 (0.01) 3.00 (0.02) 0.56 (0.21) 0.33 (0.20) 2.08 (1.27)
3000 -1.50 (0.02) 2.00 (0.01) 5.00 (0.01) 3.00 (0.01) 0.50 (0.14) 0.41 (0.15) 1.62 (0.78)
5000 -1.50 (0.01) 2.00 (0.01) 5.00 (0.00) 3.00 (0.01) 0.49 (0.12) 0.44 (0.13) 1.51 (0.62)
4.4 A Bivariate Model with a Von Mises PDF
As a final case, we use T-GIC to select the dimension of the parameter space of a model with unnormalized PDF on bounded support. Consider the bivariate von Mises PDF (Singh et al., 2002) of two circular random variables $X_{1},X_{2}$, given by
[…]
$+\lambda\sin(x_{1}-\mu_{1})\sin(x_{2}-\mu_{2})\},$ (4.8)
for $0\leq x_{1}$, $x_{2}<2\pi$, where $\kappa_{1}$, $\kappa_{2}\geq 0$, $0\leq\mu_{1}$, $\mu_{2}<2\pi$, $-\infty<\lambda<\infty$, $\theta=(\kappa_{1},\kappa_{2},\mu_{1},\mu_{2},\lambda)^{T}$ and $C(\theta)$ is the normalizing constant. The parameter $\lambda$ quantifies the dependency between two circular random variables $X_{1},X_{2}$. $C(\theta)$ is computationally intractable, involving an infinite sum of Bessel functions.
Denote $\boldsymbol{x}=(x_{1},x_{2})$ and $f_{\boldsymbol{x}}(\theta)=f(x_{1},x_{2}\mid\theta)$. Let $\{\boldsymbol{x_{i}}=(x_{i1},x_{i2}),i=1,2,\cdots,n\}$ denote a random sample from the bivariate model. After some routine calculations, we have
$\nabla_{x_{i1}}\log(f_{\boldsymbol{x_{i}}})$ $=-\kappa_{1}\sin(x_{i1}-\mu_{1})+\lambda\cos(x_{i1}-\mu_{1})\sin(x_{i2}-\mu_{2}),$ (4.9)
$\nabla_{x_{i2}}\log(f_{\boldsymbol{x_{i}}})$ $=-\kappa_{2}\sin(x_{i2}-\mu_{2})+\lambda\sin(x_{i1}-\mu_{1})\cos(x_{i2}-\mu_{2}),$ (4.10)
and
$\Delta_{x_{i1}}\log(f_{\boldsymbol{x_{i}}})$ $=-\kappa_{1}\cos(x_{i1}-\mu_{1})-\lambda\sin(x_{i1}-\mu_{1})\sin(x_{i2}-\mu_{2}),$ (4.11)
$\Delta_{x_{i2}}\log(f_{\boldsymbol{x_{i}}})$ $=-\kappa_{2}\cos(x_{i2}-\mu_{2})-\lambda\sin(x_{i1}-\mu_{1})\sin(x_{i2}-\mu_{2}).$ (4.12)
Since random variables $X_{1}$ and $X_{2}$ are independent if and only if $\lambda=0$, there are two candidate models. When $\lambda=0$, the parameter space is 4-dimensional, consisting of $(\kappa_{1},\kappa_{2},\mu_{1},\mu_{2})$. Denote this model as model $m_{1}$. For the general case with non-zero $\lambda\in R$ as defined in equation (4.8), the model is denoted as model $m_{2}$, whose parameter space is 5-dimensional. Clearly, $m_{1}$ is nested within $m_{2}$. For each model, from equations (2.2), (2.6), and (2.7), $GIC$ and its MGICE are obtained.
We conduct 100 replications to evaluate the performance of model selection between $m_{1}$ and $m_{2}$, using T-GIC1 and T-GIC2, with sample size $n=300,500,\text{and }1000$. The true parameter dimension is set to 5, with true parameter $\theta^{*}=(2.0,1.0,1.5,2.5,3.0)^{T}$. Samples are generated using the acceptance-rejection method, by generating a random number from the uniform distribution on $[0,2\pi)\times[0,2\pi)$. For numerical optimization in MGICE, we use the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm to jointly optimize all parameters, with all parameters initialized to zero.
In all replications, both T-GIC1 and T-GIC2 correctly select the true parameter dimension 5. Table 9 presents the average MGICE and its SD for parameters in the true model $m_{2}$, demonstrating the strong consistency of MGICE especially when $n$ is large.
Table 9: The average MGICE and its SD for the model $m_{2}$.
$n$ Parameter MGICE (SD)
$\kappa_{1}$ $\kappa_{2}$ $\mu_{1}$ $\mu_{2}$ $\lambda$
300 2.03 (0.20) 1.01 (0.14) 1.50 (0.05) 2.51 (0.05) 3.06 (0.24)
500 2.01 (0.14) 1.00 (0.11) 1.50 (0.04) 2.50 (0.04) 3.04 (0.17)
1000 2.01 (0.11) 1.00 (0.08) 1.50 (0.03) 2.50 (0.03) 3.01 (0.13)
5 Real Data
In this section, we apply T-GIC for model selection with real data across three domains. Now, Baker (2022) has argued that the $N\times t$ PDF provides a more realistic framework for modeling real-world data, especially in finance, and serves as an effective tool for assessing robustness and performing sensitivity analyses. In Sections 5.1 and 5.2, we consider fitting autoregression to some finance data, and polynomial regression to some car data, with $N\times t$ noise/errors, using T-GIC1 and T-GIC2 for model selection. We also compare the results with those based on AIC and BIC with Gaussian noise/errors. In Section 5.3, we consider fitting a bivariate model with von Mises PDF to some wind direction data.
5.1 Finance Data
We analyze the logged returns on the Financial Times Stock Exchange 100 Index (FTSE 100), covering 9013 observations from January 1986 to April 2021. Figure 1 shows its time plot.
Figure 1: Time plot of logged returns on the FTSE 100 from January 1986 to April 2021.
Baker (2022) assumed, for simplicity, independent data and fitted an $N\times t$ distribution. Instead, we consider an AR model to allow for data dependence. For comparison, we use T-GIC1 and T-GIC2 to select an AR model with $N\times t$ noise in equation (4.4), and AIC and BIC to select an AR model with Gaussian noise, given by
$X_{t}-c=a_{1}(X_{t-1}-c)+\cdots+a_{p}(X_{t-p}-c)+\varepsilon_{t},$ (5.1)
where $\varepsilon_{t}\sim N(0,\sigma^{2})$. Tables 10 and 11 summarize the results. Under Gaussian noise, AIC selects a seventh-order model, while BIC selects a sixth-order model. However, under $N\times t$ noise, T-GIC1 selects a sixth-order AR model, while T-GIC2 selects a second-order AR model, both yielding estimates of $\alpha$ and $k$ similar to Baker (2022). Since AR models offer a convenient framework for prediction, Table 12 presents a comparison of out-of-sample performance between the two fitted models, either chosen by AIC vs BIC or by T-GIC1 vs T-GIC2. In each case, we use the first part of the dataset for fitting and the remaining 100 data for the rolling k-step-ahead forecast. It is interesting to observe, for T-GIC1 vs T-GIC2, that the fitted AR(6) model performs better in forecasting for the short term than the fitted AR(2) model but worse for the longer term. However, for AIC vs BIC, the fitted AR(6) performs better than the fitted AR(7) uniformly.
Table 10: Selection results for AR model with two noise PDFs.
$\varepsilon_{t}\sim N(0,\sigma^{2})$
$p$ 1 2 3 4 5 6 7 8 9 10
AIC -55538.20 -55543.89 -55557.30 -55571.54 -55574.02 -55586.90 -55593.47 -55592.42 -55591.97 -55590.44
BIC -55516.88 -55515.47 -55521.77 -55528.91 -55524.27 -55530.05 -55529.51 -55521.36 -55513.80 -55505.16
$\varepsilon_{t}\sim N\times t(\alpha,k)$
$p$ 1 2 3 4 5 6 7 8 9 10
T-GIC1 12716.15 12837.14 12828.04 12812.87 12821.58 12866.07 12787.31 12862.53 12746.68 12754.23
T-GIC2 12706.13 12816.92 12797.73 12772.53 12771.13 12805.35 12716.93 12781.66 12656.55 12654.07
Table 11: Parameter estimate results for the selected AR model.
$\varepsilon_{t}\sim N(0,\sigma^{2})$
Order (method) $(a_{1},\cdots,a_{p})$ $c$ $\sigma^{2}$
$p=7$ (AIC) (-0.0004, -0.0266, -0.0448, 0.0426, -0.0215, -0.0406, 0.0308) 0.0002 0.0001
$p=6$ (BIC) (-0.0016, -0.0272, -0.0435, 0.0412, -0.0223, -0.0406) 0.0002 0.0001
$\varepsilon_{t}\sim N\times t(\alpha,k)$
Order (method) $(a_{1},\cdots,a_{p})$ $c$ s $\alpha$ $k$
$p=6$ (T-GIC1) (-0.0212, -0.0813, -0.0158, -0.0265, -0.0230, -0.0331) 0.0006 0.0129 0.0104 2.0799
$p=2$ (T-GIC2) (-0.0256, -0.0877) 0.0007 0.0129 0.0104 2.0807
Table 12: MSE and the ratio in the rolling k-step-ahead forecast on AR models.
$k$ $\varepsilon_{t}\sim N\times t(\alpha,k)$ $\varepsilon_{t}\sim N(0,\sigma^{2})$
$MSE(AR(2))$ $MSE(AR(6))$ $\frac{MSE(AR(2))}{MSE(AR(6))}$ $MSE(AR(6))$ $MSE(AR(7))$ $\frac{MSE(AR(6))}{MSE(AR(7))}$
1 $8.2026\times 10^{-5}$ $8.1961\times 10^{-5}$ 1.0008 $8.1247\times 10^{-5}$ $8.1534\times 10^{-5}$ 0.9965
2 $8.2508\times 10^{-5}$ $8.2340\times 10^{-5}$ 1.0020 $8.1492\times 10^{-5}$ $8.1763\times 10^{-5}$ 0.9967
3 $8.1973\times 10^{-5}$ $8.1565\times 10^{-5}$ 1.0050 $8.1865\times 10^{-5}$ $8.2250\times 10^{-5}$ 0.9953
4 $8.1300\times 10^{-5}$ $8.2374\times 10^{-5}$ 0.9870 $8.3796\times 10^{-5}$ $8.4129\times 10^{-5}$ 0.9960
5 $8.2529\times 10^{-5}$ $8.3208\times 10^{-5}$ 0.9918 $8.4179\times 10^{-5}$ $8.4545\times 10^{-5}$ 0.9957
5.2 Car Data
Here, we analyze the relationship between gas mileage in miles per gallon (mpg) and horsepower for 392 cars in the Auto dataset. This dataset, sourced from the StatLib library maintained at Carnegie Mellon University, was used in the 1983 American Statistical Association Exposition. Before the analysis, we apply a log transformation to mpg, as it is positive. To improve numerical stability, we also standardize horsepower. Figure 2 shows the plot of logged mpg versus standardized horsepower and the various fitted polynomial regressions.
Figure 2: Plots of logged mpg versus standardized horsepower and fits from different polynomial regression models.
The data plot suggests a nonlinear relationship between mpg and horsepower. Gareth et al. (2013) fitted a polynomial regression model with Gaussian errors to this dataset. In this study, we apply a polynomial regression model with $N\times t$ errors and use the T-GIC1 and T-GIC2 to select the degree of the polynomial, setting candidate degrees from 1 to 10. For comparison, we also use AIC and BIC with Gaussian errors. For $N\times t$ errors we employ the model in equation (4.7), and for Gaussian errors we employ the model given by
$y=\beta_{1}x+\cdots+\beta_{p}x^{p}+c+\varepsilon,$ (5.2)
where $\varepsilon\sim N(0,\sigma^{2})$.
Tables 13 and 14 report model selection and parameter estimation, respectively. Under Gaussian errors, AIC selects a seventh-degree polynomial regression model, while BIC selects a quadratic polynomial regression model. In contrast, both T-GIC1 and T-GIC2 select a quadratic polynomial regression model with $N\times t$ errors, and give estimates similar to those based on the BIC with Gaussian errors. This suggests that a quadratic model chosen by BIC is an appropriate model rather than the seventh-degree model chosen by AIC. This example shows that T-GIC1 and T-GIC2 can be used to check the choice by AIC and BIC, even when the latter can be applied, and highlights an additional utility of T-GIC coupled with the $N\times t$ PDF.
Table 13: Selection results for polynomial regression model with two error PDFs.
$\varepsilon\sim N(0,\sigma^{2})$
$p$ 1 2 3 4 5 6 7 8 9 10
AIC -186.29 -242.96 -243.44 -242.01 -249.08 -248.42 -250.44 -248.80 -248.39 -247.24
BIC -174.38 -227.07 -223.59 -218.18 -221.29 -216.65 -214.70 -209.09 -204.70 -199.59
$\varepsilon\sim N\times t(\alpha,k)$
$p$ 1 2 3 4 5 6 7 8 9 10
T-GIC1 27.69 34.24 34.07 33.89 33.08 29.71 32.66 28.84 32.38 31.93
T-GIC2 27.41 33.55 33.05 32.55 31.45 27.95 30.42 26.60 29.56 28.86
Table 14: Parameter estimate results for the selected polynomial regression model.
$\varepsilon\sim N(0,\sigma^{2})$
Degree (method) $(\beta_{1},\cdots,\beta_{p})$ $c$ $\sigma^{2}$
$p=7$ (AIC) (-0.2966, 0.0723, -0.1232, 0.0412, 0.0489, -0.0321, 0.0051) 3.0376 0.0296
$p=2$ (BIC) (-0.3448,0.0578) 3.0407 0.0309
$\varepsilon\sim N\times t(\alpha,k)$
Degree (method) $(\beta_{1},\cdots,\beta_{p})$ $c$ $s$ $\alpha$ $k$
$p=2$ (T-GIC1, T-GIC2) (-0.3838, 0.0800) 3.0288 0.3757 0.4973 3.2389
We also examine the skewness and excess kurtosis of the residuals and consider the matching between the theoretical values and their sample estimates based on the fitted residuals, as shown in Table 15. Approximate standard errors (SE) of the sample estimates are obtained by a Bootstrap procedure, with 1000 replications. For the fitted models with Gaussian errors, the skewness and the excess kurtosis of the errors are both theoretically 0. On the other hand, for $N\times t(\alpha,k)$ errors, the skewness is theoretically 0, but the excess kurtosis would not be 0 unless $k=0$. Since it is difficult to calculate the excess kurtosis for $N\times t(\alpha,k)$ PDF due to its intractable normalizing constant, we sample from $N\times t$ PDF and use the mean of the sample estimates from 1000 replications as the theoretical value, along with its SE. Apparently, the fitted residuals of the models chosen by AIC, BIC, T-GIC1 and T-GIC2 have all produced very small negative skewness of similar size, matching their theoretical value of zero reasonably well. However, for excess kurtosis, while the matching is far from being satisfactory for the Gaussian models chosen by AIC and BIC, the matching for models chosen by T-GIC1 and T-GIC2 coupled with $N\times t$ PDFs is much better.
Table 15: Comparison of residual skewness and excess kurtosis.
Noise Model Skewness Excess kurtosis
Theoretical value Sample estimate (SE) Theoretical value Sample estimate (SE)
$N(0,\sigma^{2})$ AR(7) 0 -0.08 (0.22) 0 1.13 (0.59)
AR(2) 0 -0.14 (0.20) 0 0.87 (0.49)
$N\times t(\alpha,k)$ AR(2) 0 -0.22 (0.23) 1.29 (0.21) 1.33 (0.54)
5.3 Wind Direction Data
Finally, we fit a bivariate model with von Mises PDF, as described in section 4.3, to some wind direction data, using T-GIC1 and T-GIC2 for model selection. Here, the wind direction is represented as a circular variable in radians. Matsuda et al. (2021) applied this model to wind direction data from Tokyo at 00:00 and 12:00 in 2008. For the sake of cross-validation, we analyze more recent wind direction data from Tokyo at 00:00 ($x_{1}$) and 12:00 ($x_{2}$) over 365 days in 2023, obtained from the Japan Meteorological Agency website. The data are discretized into 16 bins, such as north-northeast. Figure 3 presents the corresponding 2-d histogram.
Figure 3: 2-D histogram of wind direction data.
We fit the data using two candidate models, $m_{1}$ and $m_{2}$, representing two scenarios: $\lambda=0$ and $\lambda\in R$ in equation (4.8), respectively. For comparison, results for T-GIC1, T-GIC2 are summarized in Table 16. For $m_{2}$, both T-GIC1 and T-GIC2 are higher, indicating a better fit than $m_{1}$. This suggests that the wind directions at Tokyo on 00:00 and 12:00 in 2023 are dependent, consistent with the results in 2008 by Matsuda et al. (2021).
Table 16: Parameter estimate and T-GIC results for models $m_{1}$ and $m_{2}$.
Model Order $\kappa_{1}$ $\kappa_{2}$ $\mu_{1}$ $\mu_{2}$ $\lambda$ T-GIC1 T-GIC2
$m_{1}$ 4 0.4607 0.3057 5.6711 2.3854 0 0.1014 0.0971
$m_{2}$ 5 0.1872 0.1143 4.4243 1.3127 -1.5608 0.9857 0.9344
6 Conclusion
In this paper, based on GIC (Cheng and Tong, 2024), we propose a consistent selection criterion, T-GIC, for nested models with possibly unnormalized PDFs. Compared with existing model selection methods, T-GIC offers three advantages. First, it can handle both normalized and unnormalized PDFs. Second, it enjoys selection consistency for the case of a finite sequence of strictly nested models under mild regularity conditions. Third, it reduces significantly computational costs by avoiding the calculation of the normalizing constant and the bias correction. To showcase the efficacy of T-GIC for unnormalized PDFs, we have studied AR models and polynomial regression models with $N\times t$ noise/errors, the normalizing constants of which are typically intractable. Extensive simulation studies and real data applications have demonstrated consistency and effectiveness of T-GIC. Furthermore, we have shown excellent performance of T-GIC on PDFs with bounded support through experiments on models with bivariate von Mises PDFs and real wind direction data.
We have discussed how to use T-GIC for consistent model selection with PDFs supported on $\mathbb{R}^{d}.$ The simulation results and the real example with wind data strongly suggest that a promising direction for future research is to extend T-GIC to cover PDFs supported on bounded intervals $(a,b)$. Another direction of extension is to discrete data, thereby availing T-GIC of the opportunity of selecting an appropriate model, such as an Ising model (Friel, 2013; Everitt et al., 2017), in the area of discrete Markov random fields and spatial statistics. It is also intriguing to explore the use of combinations of different multiplying factors $C(n,k)$. Drawing inspiration from AIC and BIC, we have set $C(n,k)$ to $\exp\{-2\#(M_{k})/n\}$ and $n^{-\#(M_{k})/n}$, respectively. Although $\exp\{-2\#(M_{k})/n\}$ may not yield a consistent estimate of the true order, there is significant scope for combining it with $n^{-\#(M_{k})/n}$, similar to existing approaches for combining AIC and BIC in model selection (Hsu et al., 2019). Last, but not least, an exploration of T-GIC into non-nested models should be exciting.
Appendix A Appendix
A.1 Proof of Proposition 1
To simplify notations, we define an average operator, $P_{n}$, for data sample $x_{1},\cdots,x_{n}$ applied to any function $g(x,\cdot)$ by
$P_{n}[g(x)]=\frac{1}{n}\sum_{i=1}^{n}g(x_{i},\cdot).$
Therefore, we have that
$GIC(M(\theta))=P_{n}[W(x,p_{M(\theta)})].$
Let the constrained MGICE of $\alpha$ be $\hat{\alpha}_{0}$ while the unconstrained MGICE of $\alpha$ and $\beta$ be $\hat{\alpha}$ and $\hat{\beta}$, respectively. Note that $\hat{\alpha}_{0}$ satisfies the equation:
00 0 $=P_{n}\nabla_{\alpha}W(x,p_{M(\hat{\alpha}_{0},\beta^{*})})$
$=P_{n}\nabla_{\alpha}W(x,p_{M(\alpha^{*},\beta^{*})})+P_{n}\nabla_{\alpha}\nabla^{T}_{\alpha}W(x,p_{M(\alpha^{*},\beta^{*})})(\hat{\alpha}_{0}-\alpha^{*})+O_{p}(n^{-1})$ (A.1)
where the $O_{p}(n^{-1})$ term follows from mean value theorem and Assumption 8, and that the MGICE is root-n consistent. Similarly, the unconstrained MGICE satisfies the following equation.
$0=P_{n}\nabla_{\alpha}W(x,p_{M(\alpha^{*},\beta^{*})})+P_{n}\nabla_{\alpha}\nabla^{T}_{\alpha}W(x,p_{M(\alpha^{*},\beta^{*})})(\hat{\alpha}-\alpha^{*})$
$+P_{n}\nabla_{\alpha}\nabla^{T}_{\beta}W(x,p_{M(\alpha^{*},\beta^{*})})(\hat{\beta}-\beta^{*})+O_{p}(n^{-1}).$ (A.2)
The preceding two equations imply that
$\hat{\alpha}_{0}-\alpha^{*}=\hat{\alpha}-\alpha^{*}+D^{-1}_{\alpha^{*},\alpha^{*}}D_{\alpha^{*},\beta^{*}}(\hat{\beta}-\beta^{*})+O_{p}(n^{-1}),$ (A.3)
where $D_{a^{*},b^{*}}=P_{n}\nabla_{a}\nabla^{T}_{b}W(x,p_{M(\alpha^{*},\beta^{*})})$, with $a$, $b$ being either $\alpha$ or $\beta$, and $D_{\alpha^{*},\alpha^{*}}$ is assumed to be invertible. Doing a Taylor expansion around the constrained MGICE and after some algebra, we have
$[\log(GIC(M(\alpha^{*},\beta^{*})))-\log P_{n}\{W(x,p_{M(\hat{\alpha}_{0},\beta^{*})})\}]\times[2P_{n}\{W(x,p_{M(\alpha^{*},\beta^{*})})\}]$
$=(\hat{\alpha}_{0}-\alpha^{*})^{T}D_{\alpha^{*},\alpha^{*}}(\hat{\alpha}_{0}-\alpha^{*})+O_{p}(n^{-3/2})$
$=\begin{pmatrix}(\hat{\alpha}-\alpha^{*})^{T}&(\hat{\beta}-\beta^{*})^{T}\end{pmatrix}\begin{pmatrix}I&0\\ D_{\beta^{*},\alpha^{*}}D^{-1}_{\alpha^{*},\alpha^{*}}&0\end{pmatrix}\begin{pmatrix}D_{\alpha^{*},\alpha^{*}}&D_{\alpha^{*},\beta^{*}}\\ D_{\beta^{*},\alpha^{*}}&D_{\beta^{*},\beta^{*}}\end{pmatrix}\begin{pmatrix}I&D^{-1}_{\alpha^{*},\alpha^{*}}D_{\alpha^{*},\beta^{*}}\\ 0&0\end{pmatrix}\binom{\hat{\alpha}-\alpha^{*}}{\hat{\beta}-\beta^{*}}$
$+O_{p}(n^{-3/2})$
$=\begin{pmatrix}(\hat{\alpha}-\alpha^{*})^{T}&(\hat{\beta}-\beta^{*})^{T}\end{pmatrix}\begin{pmatrix}D_{\alpha^{*},\alpha^{*}}&D_{\alpha^{*},\beta^{*}}\\ D_{\beta^{*},\alpha^{*}}&D_{\beta^{*},\alpha^{*}}D^{-1}_{\alpha^{*},\alpha^{*}}D_{\alpha^{*},\beta^{*}}\end{pmatrix}\binom{\hat{\alpha}-\alpha^{*}}{\hat{\beta}-\beta^{*}}+O_{p}(n^{-3/2})$ (A.4)
where $I$ denotes the identity matrix of dimension $k_{0}$. Similarly, we have
$[\log(GIC(M(\alpha^{*},\beta^{*})))-\log P_{n}\{W(x,p_{M(\hat{\alpha},\hat{\beta})})\}]\times[2P_{n}\{W(x,p_{M(\alpha^{*},\beta^{*})})\}]$
$=\begin{pmatrix}(\hat{\alpha}-\alpha^{*})^{T}&(\hat{\beta}-\beta^{*})^{T}\end{pmatrix}\begin{pmatrix}D_{\alpha^{*},\alpha^{*}}&D_{\alpha^{*},\beta^{*}}\\ D_{\beta^{*},\alpha^{*}}&D_{\beta^{*},\beta^{*}}\end{pmatrix}\binom{\hat{\alpha}-\alpha^{*}}{\hat{\beta}-\beta^{*}}+O_{p}(n^{-3/2})$ (A.5)
Note that the matrix in the middle of the quadratic form depends on $n$ and it converges in probability to $D(\theta^{*})$ defined in (3.3) which is a negative-definite matrix. We can similarly partition it into a 2 by 2 block matrix:
$D(\theta^{*})=\begin{pmatrix}D(\alpha^{*},\alpha^{*})&D(\alpha^{*},\beta^{*})\\ D(\beta^{*},\alpha^{*})&D(\beta^{*},\beta^{*})\end{pmatrix}$ (A.6)
Subtracting (A.5) from (A.4) yields:
$[\log P_{n}\{W(x,p_{M(\hat{\alpha},\hat{\beta})})\}-\log P_{n}\{W(x,p_{M(\hat{\alpha}_{0},\beta^{*})})\}]\times[2P_{n}\{W(x,p_{M(\alpha^{*},\beta^{*})})\}]$
$=-(\hat{\beta}-\beta^{*})^{T}(D_{\beta^{*},\beta^{*}}-D_{\beta^{*},\alpha^{*}}D^{-1}_{\alpha^{*},\alpha^{*}}D_{\alpha^{*},\beta^{*}})(\hat{\beta}-\beta^{*})+O_{p}(n^{-3/2})$ (A.7)
However, $\log P_{n}\{W(x,p_{M(\hat{\alpha},\hat{\beta})})\}-\log P_{n}\{W(x,p_{M(\hat{\alpha}_{0},\beta^{*})})\}=\log GIC(k)-\log GIC(k_{0}).$ Recall $\sqrt{n}(\hat{\theta}-\theta^{*})$ is asymptotically normally distributed with mean zero and covariance matrix equal to $D^{-1}(\theta^{*})\Lambda(\theta^{*})D^{-T}(\theta^{*})$. It follows from routine algebra that $\sqrt{n}(\hat{\beta}-\beta^{*})$ is asymptotically normal with zero mean vector and covariance matrix equal to $B(\theta^{*})B^{T}(\theta^{*})$. Hence, $n\times[\log GIC(k)-\log GIC(k_{0})]$ converges in distribution to $Z^{T}A(\theta^{*})Z$.
A.2 Proof of Theorem 1
Proof.
First, we observe that for any model $M$ satisfying Assumptions 6-7 and assuming the validity of the law of large numbers, then it follows from routine analysis that $GIC(M(\theta))=P_{n}\{W(x,p_{M(\theta)})\}$ converges uniformly in probability to its population version $GIC_{\infty}(M(\theta))=P\{W(x,p_{M(\theta)})\}$. It follow from Assumption 7 that $GIC_{\infty}(M_{k}(\theta_{k}))$ is a Lipschitz-continuous, as a function of the parameter $\theta_{k}$, hence it attains its maximum value, denoted by $\mathcal{M}_{k}$, owing to the compact parameter space assumption (aka Assumption 6). Suppose $k_{0}$ is the smallest $k$ such that $M_{k}$ contains the true model. Then, $\mathcal{M}_{k}<\mathcal{M}_{k_{0}}$ for all $1\leq k<k_{0}$, whereas $\mathcal{M}_{k}=\mathcal{M}_{k_{0}}$ otherwise. Note that from Proposition 3 in Cheng and Tong (2024), $\mathcal{M}_{k_{0}}=H_{F}(p^{*})>0$ where $p^{*}$ is the true population pdf. Therefore, if for any $k$, $C(n,k)\to 1$ as $n\to\infty$, then the maximum T-GIC model selection criterion will not select any $k<k_{0}$, in probability.
Henceforth, consider the case that $k\geq k_{0}$. Since $\mathcal{M}_{k_{0}}>0$, $GIC(k)$ is positive, in probability, i.e., $GIC(k)>0$ holds with probability approaching 1 as sample size increases without bound. For ease of exposition, we shall assume that $GIC(k)$ is positive. Let $k>k_{0}$ be fixed. Consider the increment $D=\log\{T\text{-}GIC(k)\}-\log\{T\text{-}GIC(k_{0})\}=\log C(n,k)-\log C(n,k_{0})+\log GIC(k)-\log GIC(k_{0})$. By Proposition 1, $n\{\log GIC(k)-\log GIC(k_{0})\}$ converges weakly to some non-negative distribution. Consequently, $D=\log C(n,k)-\log C(n,k_{0})+O_{p}(1/n)$ so that if $n\times\log\{C(n,k)/C(n,k_{0})\}\to-\infty$ as $n\to\infty$, $D$ is negative in probability for $k>k_{0}$. This completes the proof of the consistency of the proposed T-GIC model selection criterion.

### v2

@@
 This method has been applied to Bayesian model selection (Dawid and Musio, 2015; Shao et al., 2019) that focuses mainly on the issue of improper priors while the data are still sampled from a normalized distribution.
-In this approach, consistent selection is only proved for non-nested models ((Shao et al., 2019), p.1826), with the situation for nested models, such as polynomial regression models or autoregressive (AR) models, remaining unclear.
+In this approach, consistent selection is only proved for non-nested models (Shao et al., 2019, p.1826), with the situation for nested models, such as polynomial regression models or autoregressive (AR) models, remaining unclear.
 Our approach is different.
@@
 For every $x\in R^{d}$ with $x=(x_{1},\cdots,x_{d})$ and for each boundary point of $x_{i},~{}i=1,\cdots,d$,
-$p(x_{1},\cdots,x_{i-1},-\infty,x_{i+1},\cdots,x_{d})\equiv 0\text{ and }p(x_{1},\cdots,x_{i-1},+\infty,x_{i+1},\cdots,x_{d})\equiv 0.$
+$p(x_{1},\cdots,x_{i-1},-\infty,x_{i+1},\cdots,x_{d})\equiv 0\text{ and }p(x_{1},\cdots,x_{i-1},+\infty,x_{i+1},\cdots,x_{d})\equiv 0,$
+where, e.g., $p(x_{1},\cdots,x_{i-1},\infty,x_{i+1},\cdots,x_{d})$ denotes $\lim_{x_{i}\to\infty}p(x_{1},\cdots,x_{d})$.
 Let $q(x)$ be a PDF on $\mathbb{R}^{d}$ satisfying Assumptions 1 and 2. The Fisher divergence between PDF $p$ and PDF $q$ is defined as
@@
 where $\Delta$ denotes the Laplacian, i.e. $\Delta_{x}f(x)=\sum_{i=1}^{d}\frac{\partial^{2}f(x)}{\partial^{2}x_{i}}$.
-Next, within their new entropy structure, Cheng and Tong (2024) define, among other notions, Gentropy $H_{G}(p)$ and G-cross-entropy $H_{G}(p,q)$ as
+Next, within their new entropy structure, Cheng and Tong (2024) defined, among other notions, Gentropy $H_{G}(p)$ and G-cross-entropy $H_{G}(p,q)$ as
 $H_{G}(p)=E_{p}||\nabla_{x}\log p(x)||^{2},$ (2.3)
@@
 Let $\theta$ be the unknown $h$-dimensional parameter vector for $p(x)$, $h\geq 1$.
-Tables 1 and 2 summarize the comparison between Shannon entropy and G-entropy across different definitions and properties.
+Tables 1 and 2 summarize a comparison between the Shannon entropy and the G-entropy (in terms of definitions and properties).
+The tables are based on the results developed by Cheng and Tong (2024).
 Table 1: Comparison between Shannon entropy and G-entropy: definitions.
@@
 Objective function $Loss(p(x))=-\log p(x)$ $W(x,p)=-||\nabla_{x}\log p(x)||^{2}~{}-~{}2\Delta_{x}\log p(x)$
-Entropy $H(x)=H(p)=E_{p}[Loss(p(x))]$ $H_{G}(x)=H_{G}(p)=E_{p}||\nabla_{x}\log p(x)||^{2}$
-Cross entropy $H(p,q)=E_{p}[Loss(q(x))]$ $H_{G}(p,q)=E_{p}[W(x,q)]$
+Entropy $H(x)=H(p)=E_{p}[-\log p(x)]$ $H_{G}(x)=H_{G}(p)=E_{p}[W(x,p)]$
+Cross entropy $H(p,q)=E_{p}[-\log q(x)]$ $H_{G}(p,q)=E_{p}[W(x,q)]$
 Joint entropy $H(x,y)=H(p_{(x,y)})$ $H_{G}(x,y)=H_{G}(p_{(x,y)})$
@@
 2.2 GIC Estimate
-Following equation (2.1), given PDF $p$, minimizing $D_{F}(p||q)$ with respect to $q$ is equivalent to maximizing the G-cross-entropy $H_{G}(p,q)$.
+By equation (2.1), given PDF $p$, minimizing $D_{F}(p||q)$ with respect to $q$ is equivalent to maximizing the G-cross-entropy $H_{G}(p,q)$.
 When $p$ is the true data PDF and $q$ the model PDF, $H_{G}(p,q)$ inspires the maximum GIC estimate (MGICE).
@@
 2.3 GIC for Model Selection
-Consider a collection of candidate parametric models $M_{1},\cdots,M_{K}$, denoted as $M_{k}(\theta_{k}),k=1,\cdots,K$, with $\theta_{k}$ ($h_{k}\geq 1$) being an unknown $h_{k}$-dimensional parameter vector in parameter space $\Theta.$ Similar to AIC, an unbiased $GIC_{c}$ for model selection (Cheng and Tong 2024) is derived under mild regularity conditions by correcting the bias, $B$, in $nGIC$ as follows:
+Consider a collection of candidate parametric models $M_{1},\cdots,M_{K}$, denoted as $M_{k}(\theta_{k}),k=1,\cdots,K$, with $\theta_{k}$ ($h_{k}\geq 1$) being an unknown $h_{k}$-dimensional parameter vector in parameter space $\Theta.$ Similar to AIC, an unbiased $GIC_{c}$ for model selection (Cheng and Tong 2024) is derived under mild regularity conditions by correcting the bias, $B$, in $n\times GIC$ as follows:
 $GIC_{c}(M_{k}(\hat{\theta}_{k}))=GIC_{n}(M_{k}(\hat{\theta}_{k}))-B_{k},$ (2.8)
-where $\hat{\theta}_{k}$ is the MGICE of $\theta_{k}$, $GIC_{n}(M_{k}(\hat{\theta}_{k}))=nGIC(M_{k}(\hat{\theta}_{k}))$, and
+where $\hat{\theta}_{k}$ is the MGICE of $\theta_{k}$, $GIC_{n}(M_{k}(\hat{\theta}_{k}))=n\times GIC(M_{k}(\hat{\theta}_{k}))$, and
 $B_{k}$ $=-tr\{E_{p_{x}}[\nabla_{\theta}W(x,p_{M_{i}(\theta^{*})})\nabla_{\theta}^{T}W(x,p_{M_{i}(\theta^{*})})]E_{p_{x}}^{-1}[\nabla_{\theta}^{2}W(x,p_{M_{i}(\theta^{*})})]\}.$ (2.9)
@@
 The $T\text{-}GIC(k)$ of model $M_{k}$ is defined as
-$T\text{-}GIC(k)=C(n,k)GIC(M_{k}(\hat{\theta}_{k})),$ (3.1)
+$T\text{-}GIC(k)=C(n,k)\times GIC(M_{k}(\hat{\theta}_{k})),$ (3.1)
 where $C(n,k)$ is a constant depending only on $n$ and $k$, $GIC(M_{k}(\hat{\theta}_{k}))$ is defined by equation (2.6), and $\hat{\theta}_{k}$ is the MGICE of $\theta_{k}$ for model $M_{k}$.
@@
 Assumption 7.
-Both $\nabla_{x}^{2}\log p_{M(\theta)}(x)$ and $[\nabla_{x}\log p_{M(\theta)}(x)][\nabla_{x}\log p_{M(\theta)}(x)]^{T}$ are Lipschitz continuous in respect of Frobenious norm.
+Both $\nabla_{x}^{2}\log p_{M(\theta)}(x)$ and $[\nabla_{x}\log p_{M(\theta)}(x)][\nabla_{x}\log p_{M(\theta)}(x)]^{T}$ are Lipschitz continuous in respect of Frobenius norm.
 Specifically, $\forall\theta_{1},\theta_{2}\in\Theta$,
@@
 Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model.
-If $K=k_{0}$, under Assumptions 1-8, then $T\textbf{-}GIC1=\exp\{-2\#(M_{k})/n\}GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
+If $K=k_{0}$, under Assumptions 1-8, then $T\textbf{-}GIC1=\exp\{-2\#(M_{k})/n\}\times GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
 Corollary 2.
 Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model.
-Under Assumptions 1-8, then $T\textbf{-}GIC2=n^{-\#(M_{k})/n}GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
+Under Assumptions 1-8, then $T\textbf{-}GIC2=n^{-\#(M_{k})/n}\times GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
 However, similar to AIC, T-GIC1 may fail to provide consistent selection when $K>k_{0}$.
@@
 Samples are generated using the acceptance-rejection method, by generating random numbers from the $N(0,1/\alpha)$ distribution.
-For numerical optimization in MGICE, we use the Adaptive Moment Estimation (Adam) algorithm to jointly optimize all parameters.
+For numerical optimization in MGICE, we use the Adaptive Moment Estimation (Adam) algorithm (Kingma and Ba, 2015) to jointly optimize all parameters.
 Similar to Baker (2022), for all experiments involving the $N\times t$ PDF, we use the sample mean and standard deviation as starting values for $\mu$ and $s$, and regular initial values for $\alpha$ and $k$, e.g.
@@
 We conduct 100 replications to obtain the frequency distribution of the selected model orders, ranging from 1 to 10, using T-GIC1 and T-GIC2, with sample size $N=1000,3000,$ and $5000$.
-The true model order $p^{*}$ is set to 3, with true parameters $(a_{1}^{*},a_{2}^{*},a_{3}^{*},c,s,\alpha,k)$ specified at $(0.50,-0.25,0.10,3.00,0.50,0.50,1.50)$.
+The true model order $p^{*}$ is set to 3, with true parameters $(a_{1}^{*},a_{2}^{*},a_{3}^{*},c^{*},s^{*},\alpha^{*},k^{*})$ specified at $(0.50,-0.25,0.10,3.00,0.50,0.50,1.50)$.
 In each replication, the first 200 data are discarded to ensure stationarity.
@@
 We conduct 100 replications to obtain the frequency distribution of the selected model degree $p$, ranging from 1 to 10, using T-GIC1 and T-GIC2, with sample size $n=1000,3000,$ and $5000$.
-The true model order $p^{*}$ is set to 3, with true parameters $(\beta_{1}^{*},\beta_{2}^{*},\beta_{3}^{*},c,s,\alpha,k)$ specified as $(-1.5,2.0,5.0,3.0,0.5,0.5,1.5)$.
+The true model order $p^{*}$ is set to 3, with true parameters $(\beta_{1}^{*},\beta_{2}^{*},\beta_{3}^{*},c^{*},s^{*},\alpha^{*},k^{*})$ specified as $(-1.5,2.0,5.0,3.0,0.5,0.5,1.5)$.
 Table 7 suggests that both T-GIC1 and T-GIC2 consistently select the correct degree in probability as $n$ increases.
@@
 Samples are generated using the acceptance-rejection method, by generating a random number from the uniform distribution on $[0,2\pi)\times[0,2\pi)$.
-For numerical optimization in MGICE, we use the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm to jointly optimize all parameters, with all parameters initialized to zero.
+For numerical optimization in MGICE, we use the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm (Nocedal and Wright, 2006) to jointly optimize all parameters, with all parameters initialized to zero.
 In all replications, both T-GIC1 and T-GIC2 correctly select the true parameter dimension 5.
@@
 Since AR models offer a convenient framework for prediction, Table 12 presents a comparison of out-of-sample performance between the two fitted models, either chosen by AIC vs BIC or by T-GIC1 vs T-GIC2.
-In each case, we use the first part of the dataset for fitting and the remaining 100 data for the rolling k-step-ahead forecast.
+In each case, we use the first part of the dataset for fitting and the remaining 100 data for the rolling $m$-step-ahead forecast.
 It is interesting to observe, for T-GIC1 vs T-GIC2, that the fitted AR(6) model performs better in forecasting for the short term than the fitted AR(2) model but worse for the longer term.
 However, for AIC vs BIC, the fitted AR(6) performs better than the fitted AR(7) uniformly.
+Overall, Table 12 suggests that the effect of heavy-tailed innovation kicks in for the longer term forecasting.
 Table 10: Selection results for AR model with two noise PDFs.
@@
 $p=2$ (T-GIC2) (-0.0256, -0.0877) 0.0007 0.0129 0.0104 2.0807
-Table 12: MSE and the ratio in the rolling k-step-ahead forecast on AR models.
-$k$ $\varepsilon_{t}\sim N\times t(\alpha,k)$ $\varepsilon_{t}\sim N(0,\sigma^{2})$
+Table 12: MSE and the ratio in the rolling m-step-ahead forecast on AR models. (Row-wise minimum values are underlined.)
+$m$ $\varepsilon_{t}\sim N\times t(\alpha,k)$ $\varepsilon_{t}\sim N(0,\sigma^{2})$
 $MSE(AR(2))$ $MSE(AR(6))$ $\frac{MSE(AR(2))}{MSE(AR(6))}$ $MSE(AR(6))$ $MSE(AR(7))$ $\frac{MSE(AR(6))}{MSE(AR(7))}$
@@
 Then, $\mathcal{M}_{k}<\mathcal{M}_{k_{0}}$ for all $1\leq k<k_{0}$, whereas $\mathcal{M}_{k}=\mathcal{M}_{k_{0}}$ otherwise.
-Note that from Proposition 3 in Cheng and Tong (2024), $\mathcal{M}_{k_{0}}=H_{F}(p^{*})>0$ where $p^{*}$ is the true population pdf.
+Note that from Proposition 3 in Cheng and Tong (2024), $\mathcal{M}_{k_{0}}=H_{G}(p^{*})>0$ where $p^{*}$ is the true population pdf.
 Therefore, if for any $k$, $C(n,k)\to 1$ as $n\to\infty$, then the maximum T-GIC model selection criterion will not select any $k<k_{0}$, in probability.
@@
 Consider the increment $D=\log\{T\text{-}GIC(k)\}-\log\{T\text{-}GIC(k_{0})\}=\log C(n,k)-\log C(n,k_{0})+\log GIC(k)-\log GIC(k_{0})$.
-By Proposition 1, $n\{\log GIC(k)-\log GIC(k_{0})\}$ converges weakly to some non-negative distribution.
+By Proposition 1, $n\times\{\log GIC(k)-\log GIC(k_{0})\}$ converges weakly to some non-negative distribution.
 Consequently, $D=\log C(n,k)-\log C(n,k_{0})+O_{p}(1/n)$ so that if $n\times\log\{C(n,k)/C(n,k_{0})\}\to-\infty$ as $n\to\infty$, $D$ is negative in probability for $k>k_{0}$.

### v3

@@
-On a fast consistent selection of nested models with possibly unnormalised probability densities
+On a fast consistent selection of nested models with possibly unnormalized probability densities
 Abstract
@@
 Existing methods to address this issue often incur high computational costs, either due to numerical approximations of normalizing constants or evaluation of bias corrections in information criteria.
-In this paper, we propose a novel and fast selection criterion, T-GIC, for nested models, allowing direct data sampling from a possibly unnormalized probability density function.
-T-GIC gives a consistent selection under mild regularity conditions and is computationally efficient, benefiting from a multiplying factor that depends only on the sample size and the model complexity.
-Extensive simulation studies and real-data applications demonstrate the efficacy of T-GIC in the selection of nested models with unnormalized probability densities.
-Keywords: unnormalized probability densities, gradient-based information criterion, consistent model selection, computational efficiency, nested models
+In this paper, we propose a novel and fast selection criterion, MIC, for nested models of possibly dependent data, allowing direct data sampling from a possibly unnormalized probability density function.
+With a suitable multiplying factor depending only on the sample size and the model complexity, MIC gives a consistent selection under mild regularity conditions and is computationally efficient.
+Extensive simulation studies and real-data applications demonstrate the efficacy of MIC in the selection of nested models with unnormalized probability densities.
+Keywords: Unnormalized probability densities, Gradient-based information criterion, Consistent model selection, Computational efficiency, Nested models, Markov dependent data
 1 Introduction
 Models with unnormalized probability density functions are ubiquitous.
-In statistics, artificial intelligence, statistical mechanics and many other fields, often we only want or are able to stipulate the general shape of the models’ distributions without requiring their underlying probability density functions (PDFs) to integrate to unity, hence the notion of unnormalized PDFs.
-In such cases, the normalizing constants are either extremely difficult or impossible to compute directly.
+In statistics, artificial intelligence, statistical mechanics and many other fields, often we only want or are able to stipulate the general shape of the models’ distributions without requiring the underlying probability density functions (PDFs) to integrate to unity, hence the notion of unnormalized PDFs.
+Here, the normalizing constants are either difficult or impossible to compute explicitly.
 They arise in various circumstances, e.g.
-non-conjugacy in Bayesian posteriors (Andrade and Rathie, 2017), partition functions in statistical mechanics (Frigg and Werndl, 2024), directional distributions of data on a sphere (Pewsey and García-Portugués, 2021), Ising model in spatial statistics (Friel, 2013), and distributions with high-dimensional latent variables (Murray and Salakhutdinov, 2008).
+non-conjugacy in Bayesian posteriors (Andrade & Rathie 2017), partition functions in statistical mechanics (Frigg & Werndl 2024), directional distributions of data on a sphere (Pewsey & García-Portugués 2021), Ising model in spatial statistics (Friel 2013), and distributions with high-dimensional latent variables (Murray & Salakhutdinov 2008).
 In short, unnormalized PDFs pose a significant challenge in likelihood-based model comparison.
-Addressing the challenge, we focus on the selection of nested models with possibly unnormalized PDFs.
-Model selection involves choosing the best statistical model from several candidates based on the observed data, with each candidate having possibly a different number of parameters (Rao et al., 2001).
-Among the forerunners and the most prominent members of this vast industry, Akaike (1974) proposed the Akaike Information Criterion (AIC) based on the Kullback-Leibler divergence that is underpinned by the Shannon entropy.
-Here, the model with the minimum AIC is selected as the best model.
+Addressing it, we focus on the selection of nested models with possibly unnormalized PDFs.
+Model selection involves choosing the best statistical model from several candidates based on the observed data, with each candidate having possibly a different number of parameters (Rao et al.
+2001).
+Among the forerunners, Akaike (1974) proposed the Akaike Information Criterion (AIC) based on the Kullback-Leibler divergence.
 As the other, Schwarz (1978) developed the Bayesian Information Criterion (BIC) via a Laplace approximation of the Bayes factor.
-Here, the model with the minimum BIC value is selected.
 Unfortunately, none of them is applicable for the selection of models with unnormalized PDFs.
-At present, there are essentially two main approaches to overcome the challenge.
-One approach is to first estimate the intractable normalizing constant by numerical approximations or Markov Chain Monte Carlo.
-See, for example, (Baker, 2022) and (MacKay, 2003; Congdon, 2006).
+There are two main approaches to overcome the challenge.
+One is to first estimate the intractable normalizing constant by numerical approximations or Markov Chain Monte Carlo.
+See, for example, (Baker 2022) and (MacKay 2003, Congdon 2006).
 While simple and intuitive, this approach is computationally intensive, especially for random vectors in high-dimensional settings.
-Another approach is to avoid the calculation of the normalizing constant altogether.
-For this, the score matching method (Hyvärinen, 2005) provides significant potential through efficient computation, as it does not depend on the normalizing constant.
-This method has been applied to Bayesian model selection (Dawid and Musio, 2015; Shao et al., 2019) that focuses mainly on the issue of improper priors while the data are still sampled from a normalized distribution.
-In this approach, consistent selection is only proved for non-nested models (Shao et al., 2019, p.1826), with the situation for nested models, such as polynomial regression models or autoregressive (AR) models, remaining unclear.
-Our approach is different.
-We allow the data to be sampled directly from an unnormalized PDF.
-Recently, Matsuda et al.
+Another is to avoid the calculation of the normalizing constant altogether via the score matching method (Hyvärinen 2005) that provides significant potential through efficient computation without the normalizing constant.
+This method has been applied to Bayesian model selection (Dawid & Musio 2015, Shao et al.
+2019) that focuses mainly on the issue of improper priors while the data are still sampled from a normalized distribution.
+In this approach, consistent selection is only proved for non-nested models (Shao et al.
+2019, p.1826), with the situation for nested models, such as polynomial regression models or autoregressive (AR) models, remaining unclear.
+Direct sampling from an unnormalized PDF in model selection is challenging.
+For this, the following approaches are available.
+Recently, for independent data, Matsuda et al.
 (2021) proposed an information criterion for selection of models with unnormalized PDFs estimated via noise contrastive estimation (NCE) or score matching.
-Slightly later, Cheng and Tong (2024) introduced a new entropy, the G-entropy, that includes an information criterion for model selection via a negative Hyvärinen score.
+Slightly later, for both independent and Markov dependent data, Cheng & Tong (2024) proposed an information criterion for model selection.
 In the above information criteria, bias correction is involved and the estimation of the bias term can be computationally intensive.
-For example, with parameter dimension $k$ and sample size $n$, matrix calculations for the bias in Cheng and Tong (2024) run in $O(nk^{2}+k^{3})$ time.
+For example, with parameter dimension $k$ and sample size $n$, matrix calculations for the bias in Cheng & Tong (2024) run in $\mathcal{O}(nk^{2}+k^{3})$ time.
 Similar order applies to Matsuda et al.
-(2021)
-In this paper, we propose a fast information criterion for consistent selection of nested models with unnormalized PDFs.
-We name the criterion T-GIC.
-In it, we introduce a multiplying factor $C(n,k)$, which depends only on the sample size $n$ and the order $k$ of the candidate model $M_{k}$, resulting in a drastic reduction of computation to $O(1)$.
-We show that T-GIC achieves consistent model selection for strictly nested models under mild regularity conditions.
-Further, we demonstrate the efficacy of T-GIC through simulation on AR and polynomial regression models with unnormalized PDFs as well as a model with a bivariate von Mises PDF with bounded support.
-Finally, we apply the T-GIC to real data from diverse domains, including finance, automotive engineering, and wind direction analysis.
-This paper is organized as follows.
-In Section 2, we provide a brief review of F-divergence and associated notions.
-In Section 3, we introduce the T-GIC and prove its consistency for model selection under regularity assumptions.
-In Section 4, we conduct simulations to demonstrate the efficacy of T-GIC.
-In Section 5, we apply T-GIC to various real-world data, offering some insights.
-In Section 6, we conclude with a discussion of our findings and potential directions for future research.
-Proofs are provided in Appendix A.
+(2021).
+In this paper, we propose a fast information criterion for consistent selection of nested models of possibly Markov dependent data with unnormalized PDFs.
+We name the criterion MIC.
+In it, we introduce a multiplying factor $C(n,k)$, which depends only on the sample size $n$ and the order $k$ of the candidate model $M_{k}$, resulting in a drastic reduction of computation to $\mathcal{O}(1)$.
+We show that MIC achieves consistent model selection for strictly nested models under mild regularity conditions.
+Further, we demonstrate the efficacy of MIC through simulation on AR and polynomial regression models with unnormalized PDFs as well as a model with a bivariate von Mises PDF with bounded support.
+Finally, we apply the MIC to real data from diverse domains, including finance, automotive engineering, and wind direction analysis.
+[…]
 2 A New Information Criterion
-It is well known that the Kullback-Leibler divergence is underpinned by the Shannon entropy.
-Recently, Cheng and Tong (2024) have developed a new entropy that parallels Shannon’s.
-It is called the G-entropy (sometimes shortened to Gentropy).
-They have shown how it underpins the Fisher divergence and leads to the GIC for model selection.
-2.1 Fisher Divergence and Gentropy
-Let $p(x)$ be a PDF on $\mathbb{R}^{d}$ under the following assumptions:
+2.1 Fisher Divergence
+Let $p(x)$ be a generic PDF on $\mathbb{R}^{d}$ under the following assumptions:
 Assumption 1.
@@
 For every $x\in R^{d}$ with $x=(x_{1},\cdots,x_{d})$ and for each boundary point of $x_{i},~i=1,\cdots,d$,
-$p(x_{1},\cdots,x_{i-1},-\infty,x_{i+1},\cdots,x_{d})\equiv 0\text{ and }p(x_{1},\cdots,x_{i-1},+\infty,x_{i+1},\cdots,x_{d})\equiv 0,$
+[…]
 where, e.g., $p(x_{1},\cdots,x_{i-1},\infty,x_{i+1},\cdots,x_{d})$ denotes $\lim_{x_{i}\to\infty}p(x_{1},\cdots,x_{d})$.
-Let $q(x)$ be a PDF on $\mathbb{R}^{d}$ satisfying Assumptions 1 and 2. The Fisher divergence between PDF $p$ and PDF $q$ is defined as
-$D_{F}(p||q)=E_{p}||\nabla_{x}\log p(x)~{}-~{}\nabla_{x}\log q(x)||^{2}.$ (2.1)
-The objective function $W$ of PDF $q$ at point $x$ is a negative Hyvärinen score (Hyvärinen, 2005), defined by
-$W(x,q)=-||\nabla_{x}\log q(x)||^{2}~{}-~{}2\Delta_{x}\log q(x),$ (2.2)
-where $\Delta$ denotes the Laplacian, i.e. $\Delta_{x}f(x)=\sum_{i=1}^{d}\frac{\partial^{2}f(x)}{\partial^{2}x_{i}}$.
-Next, within their new entropy structure, Cheng and Tong (2024) defined, among other notions, Gentropy $H_{G}(p)$ and G-cross-entropy $H_{G}(p,q)$ as
-$H_{G}(p)=E_{p}||\nabla_{x}\log p(x)||^{2},$ (2.3)
-$H_{G}(p,q)=E_{p}[W(x,q)].$ (2.4)
-Integration by parts yields the following relationship
-$D_{F}(p||q)=H_{G}(p)-H_{G}(p,q).$
-Denote $x$ and $y$ as two random vectors with joint PDF $p(x,y)$.
+Consider the following objective function:
+[…]
+where $\Delta$ denotes the Laplacian, i.e.
+$\Delta_{x}f(x)=\sum_{i=1}^{d}\frac{\partial^{2}f(x)}{\partial^{2}x_{i}}$.
+For PDFs $p$ and $q$ under the above assumptions, the Fisher divergence satisfies the following equation:
+$D_{F}(p||q)=E_{p}||\nabla_{x}\log p(x)~-~\nabla_{x}\log q(x)||^{2}=E_{p}[W(x,p)]-E_{p}[W(x,q)].$ (2.2)
+See, e.g., Cheng & Tong (2024).
+Note that $E_{p}[W(x,p)]$, interpreted as an entropy, is $trace(G_{p}),$ where $G_{p}$ is a matrix whose $(i,j)$-th element is $E_{p}[f_{i}(x)f_{j}(x)/p^{2}(x)],$ $f_{i}(x)=\partial p/\partial x_{i}.$ See (Cheng & Tong (to appear in 2025)), who introduced the matrix under the name covariate Fisher Information Matrix based on a Riemannian geometry approach.
+2.2 GIC Estimate
 Let $\theta$ be the unknown $h$-dimensional parameter vector for $p(x)$, $h\geq 1$.
-Tables 1 and 2 summarize a comparison between the Shannon entropy and the G-entropy (in terms of definitions and properties).
-The tables are based on the results developed by Cheng and Tong (2024).
-Table 1: Comparison between Shannon entropy and G-entropy: definitions.
-Definitions Shannon entropy G-entropy
-Objective function $Loss(p(x))=-\log p(x)$ $W(x,p)=-||\nabla_{x}\log p(x)||^{2}~{}-~{}2\Delta_{x}\log p(x)$
-Entropy $H(x)=H(p)=E_{p}[-\log p(x)]$ $H_{G}(x)=H_{G}(p)=E_{p}[W(x,p)]$
-Cross entropy $H(p,q)=E_{p}[-\log q(x)]$ $H_{G}(p,q)=E_{p}[W(x,q)]$
-Joint entropy $H(x,y)=H(p_{(x,y)})$ $H_{G}(x,y)=H_{G}(p_{(x,y)})$
-Conditional entropy $H(y|x)=E_{p_{(x,y)}}[Loss(\frac{p(x,y)}{p(x)})]$ $H_{G}(y|x)=E_{p(x,y)}[W((x,y),\frac{p(x,y)}{p(x)})]$
-Mutual information $I(x,y)=H(x)+H(y)-H(x,y)$ $I_{G}(x,y)=H_{G}(x,y)-H_{G}(x)-H_{G}(y)$
-Information matrix $IM_{\theta}=E_{p}[[\nabla_{\theta}\log p(x)][\nabla_{\theta}\log p(x)]^{\tau}]$ $GIM_{x}=E_{p}[[\nabla_{x}\log p(x)][\nabla_{x}\log p(x)]^{\tau}]$
-Divergence $D_{KL}(p||q)=E_{p}[Loss(\frac{q(x)}{p(x)})]$ $=H(p,q)-H(p)$ $D_{F}(p||q)=E_{p}||\nabla_{x}\log p(x)~{}-~{}\nabla_{x}\log q(x)||^{2}$ $=H_{G}(p)-H_{G}(p,q)$
-Table 2: Comparison between Shannon entropy and G-entropy: properties.
-Shannon entropy G-entropy
-1 $H(p,p)=H(p)=E_{p}[Loss(p(x))]$ $H_{G}(p,p)=H_{G}(p)=E_{p}[W(x,p)]$
-2 $H(p)\leq H(p,q)$ $H_{G}(p)\geq H_{G}(p,q)$
-3 $\min_{q}\{H(p,q)\}=H(p)$ $\max_{q}\{H_{G}(p,q)\}=H_{G}(p)$
-4 $H(x,y)\leq H(x)+H(y)$, with equality iff $x$ and $y$ are independent. $H_{G}(x,y)\geq H_{G}(x)+H_{G}(y)$, with equality iff $x$ and $y$ are independent.
-5 $H(y|x)=H(x,y)-H(x)$ $H_{G}(y|x)=H_{G}(x,y)-H_{G}(x)$
-2.2 GIC Estimate
-By equation (2.1), given PDF $p$, minimizing $D_{F}(p||q)$ with respect to $q$ is equivalent to maximizing the G-cross-entropy $H_{G}(p,q)$.
-When $p$ is the true data PDF and $q$ the model PDF, $H_{G}(p,q)$ inspires the maximum GIC estimate (MGICE).
+When $p$ is the true data PDF and $q$ the model PDF, by equation (2.2), given PDF $p$ minimizing $D_{F}(p||q)$ with respect to $q$ is equivalent to maximizing $E_{p}[W(x,q)],$ inspiring the maximum GIC estimate (MGICE).
 Specifically, let $p_{x}$ be the data PDF, and $p_{M(\theta)}$ be a PDF of model $M$ with unknown $h$-dimensional parameter vector $\theta$ ($h\geq 1$) in parameter space $\Theta$.
 The true parameter $\theta^{*}$ can be obtained by
-$\theta^{*}=\arg\min_{\theta\in\Theta}D_{F}(p_{x}||p_{M(\theta)})=\arg\max_{\theta\in\Theta}H_{G}(p_{x},p_{M(\theta)}).$ (2.5)
-Let $x_{1},\cdots,x_{n}$ be a sample in $\mathbb{R}^{d}$ from the data PDF $p_{x}$. An unbiased estimate of $H_{G}(p_{x},p_{M(\theta)})$ is given by
+$\theta^{*}=\arg\min_{\theta\in\Theta}D_{F}(p_{x}||p_{M(\theta)})=\arg\max_{\theta\in\Theta}E_{p}[W(x,p_{M(\theta)})].$ (2.3)
+Let $x_{1},\cdots,x_{n}$ be a sample in $\mathbb{R}^{d}$ from the data PDF $p_{x}$. An unbiased estimate of $E_{p}[W(x,p_{M(\theta)})]$ is given by
 $GIC(M(\theta))=\frac{1}{n}\sum_{i=1}^{n}W(x_{i},p_{M(\theta)}).$ (2.4)
@@
 2.3 GIC for Model Selection
-Consider a collection of candidate parametric models $M_{1},\cdots,M_{K}$, denoted as $M_{k}(\theta_{k}),k=1,\cdots,K$, with $\theta_{k}$ ($h_{k}\geq 1$) being an unknown $h_{k}$-dimensional parameter vector in parameter space $\Theta.$ Similar to AIC, an unbiased $GIC_{c}$ for model selection (Cheng and Tong 2024) is derived under mild regularity conditions by correcting the bias, $B$, in $n\times GIC$ as follows:
+Consider a collection of candidate parametric models $M_{1},\cdots,M_{K}$, denoted as $M_{k}(\theta_{k}),k=1,\cdots,K$, with $\theta_{k}$ ($h_{k}\geq 1$) being an unknown $h_{k}$-dimensional parameter vector in parameter space $\Theta.$ An unbiased $GIC_{c}$ for model selection is derived under mild regularity conditions by correcting the bias, $B$, in $n\times GIC$ as follows:
 $GIC_{c}(M_{k}(\hat{\theta}_{k}))=GIC_{n}(M_{k}(\hat{\theta}_{k}))-B_{k},$ (2.6)
@@
 By maximizing $GIC_{c}$, an appropriate model from $M_{1},\cdots,M_{K}$ is selected.
-3 A New Model Selection Criterion: T-GIC
-In this section, we propose a fast model selection criterion, T-GIC, and show the consistency for a finite sequence of strictly nested models under mild regularity conditions.
+3 A New Model Selection Criterion: MIC
+In this section, we propose a fast model selection criterion, MIC, and show the consistency for a finite sequence of strictly nested models under mild regularity conditions.
 3.1 A Fast Model Selection
@@
 Suppose we have a collection of candidate parametric models as described in Section 2.3.
-The $T\text{-}GIC(k)$ of model $M_{k}$ is defined as
-$T\text{-}GIC(k)=C(n,k)\times GIC(M_{k}(\hat{\theta}_{k})),$ (3.1)
+The $MIC(k)$ of model $M_{k}$ is defined as
+$MIC(k)=C(n,k)\times GIC(M_{k}(\hat{\theta}_{k})),$ (3.1)
 where $C(n,k)$ is a constant depending only on $n$ and $k$, $GIC(M_{k}(\hat{\theta}_{k}))$ is defined by equation (2.4), and $\hat{\theta}_{k}$ is the MGICE of $\theta_{k}$ for model $M_{k}$.
-We propose to select the model that maximizes $T\text{-}GIC(k)$.
-Compared to $GIC_{c}$ in equation (2.8), $T\text{-}GIC$ significantly reduces the computational costs by introducing a factor $C(n,k)$ to bypass the bias correction calculation.
-Note that estimating the bias, B, based on a sample of size $n$ involves a calculation $O(nh_{k}^{2}+h_{k}^{3})$, which becomes computationally expensive when $h_{k}$ or $n$ is large.
-In contrast, using the multiplying factor $C(n,k)$ reduces the computational cost to only $O(1)$.
-The following section shows the consistency of the proposed model selection criterion T-GIC.
-3.2 Consistency of T-GIC
-We first introduce several assumptions, which have been used by Song et al. (2020) and Cheng and Tong (2024).
+We propose to select the model that maximizes $MIC(k)$.
+The following sections demonstrate the high computational efficiency and the selection consistency of the proposed criterion $MIC$.
+3.2 Computational cost comparison
+[…]
+Table 1: Computational cost of penalty terms in model selection criteria.
+Criteria GICc NCIC1 NCIC2 MIC
+[…]
+3.3 Consistency of MIC
+The following assumptions are from Song et al. (2020) and Cheng & Tong (2024).
 Assumption 4.
@@
 Assumption 7.
-Both $\nabla_{x}^{2}\log p_{M(\theta)}(x)$ and $[\nabla_{x}\log p_{M(\theta)}(x)][\nabla_{x}\log p_{M(\theta)}(x)]^{T}$ are Lipschitz continuous in respect of Frobenius norm.
+Both $\nabla_{x}^{2}\log p_{M(\theta)}(x)$ and $\nabla_{x}\log p_{M(\theta)}(x)][\nabla_{x}\log p_{M(\theta)}(x)]^{T}$ are Lipschitz continuous in respect of Frobenius norm.
 Specifically, $\forall\theta_{1},\theta_{2}\in\Theta$,
-$||\nabla_{x}^{2}\log p_{M(\theta_{1})}(x)-\nabla_{x}^{2}\log p_{M(\theta_{2})}(x)||_{F}\leq L_{1}(x)||\theta_{1}-\theta_{2}||_{2},$
+[…]
 and
-$||[\nabla_{x}\log p_{M(\theta_{1})}(x)][\nabla_{x}\log p_{M(\theta_{1})}(x)]^{T}-[\nabla_{x}\log p_{M(\theta_{2})}(x)][\nabla_{x}\log p_{M(\theta_{2})}(x)]^{T}||$
-$\leq L_{2}(x)||\theta_{1}-\theta_{2}||_{2}.$
+$\nabla_{x}\log p_{M(\theta_{1})}(x)][\nabla_{x}\log p_{M(\theta_{1})}(x)]^{T}-[\nabla_{x}\log p_{M(\theta_{2})}(x)][\nabla_{x}\log p_{M(\theta_{2})}(x)]^{T}||$
+[…]
 In addition, $E_{p_{x}}[L_{1}^{2}(x)]<\infty$ and $E_{p_{x}}[L_{2}^{2}(x)]<\infty.$
@@
 For $\theta_{1}$, $\theta_{2}$ near $\theta^{*}$, and $\forall i,j$,
-$||\nabla_{\theta}^{2}\partial_{i}\partial_{j}\log p_{M(\theta_{1})}-\nabla_{\theta}^{2}\partial_{i}\partial_{j}\log p_{M(\theta_{2})}||_{F}\leq M_{i,j}(x)||\theta_{1}-\theta_{2}||_{2},$
+[…]
 and
-$||\nabla_{\theta}^{2}\partial_{i}\log p_{M(\theta_{1})}\partial_{j}\log p_{M(\theta_{1})}-\nabla_{\theta}^{2}\partial_{i}\log p_{M(\theta_{1})}\partial_{j}\log p_{M(\theta_{2})}||_{F}\leq N_{i,j}(x)||\theta_{1}-\theta_{2}||_{2}.$
+[…]
 Here, $\partial_{i}$ refers to the partial derivative with respect to the component $x_{i}$ in the random vector $x=(x_{1},\cdots,x_{d}).$
@@
 Assumption 7 defines Lipschitz continuity, while Assumption 8 describes Lipschitz smoothness for second derivatives.
-Based on these assumptions, Proposition 6 in Cheng and Tong (2024) shows the asymptotic normality of the MGICE, as stated in Lemma 1.
+Based on these assumptions, Proposition 6 in Cheng & Tong (2024) shows the asymptotic normality of the MGICE, as stated in Lemma 1.
 Lemma 1.
@@
 Aspects MLE MGICE
-Objective function $\log[p_{M(\theta)}(x_{i})]$ $W(x_{i},p_{M(\theta)})=-||\nabla_{x}\log p_{M(\theta)}(x_{i})||^{2}~{}-~{}2\Delta_{x}\log p_{M(\theta)}(x_{i})$
+[…]
 Sample function $l_{n}(M(\theta))=\sum_{i=1}^{n}\log[p_{M(\theta)}(x_{i})]$ $GIC_{n}(M(\theta))=\sum_{i=1}^{n}W(x_{i},p_{M(\theta)})$
@@
 where $Z$ is a $(k-k_{0})$-dimensional standard normal random vector, $p^{*}=p_{M(\alpha^{*},\beta^{*})}$, and $I$ denotes the identity matrix of dimension $k_{0}$.
-Now, we establish the consistency of our T-GIC method.
+Now, we establish the consistency of our MIC method.
 Theorem 1.
@@
     $C(n,k)$ is such that (i) for any $k$, $C(n,k)\to 1$ as $n\to\infty$ and (ii) for any $K\geq k_{1}>k_{2}\geq 1$, $n\times\log\{C(n,k_{1})/C(n,k_{2})\}\to$-$\infty$ as $n\to\infty$.
-Let $\hat{k}=\arg\max_{1\leq k\leq K}T\text{-}GIC(k)$. Then $\hat{k}$ converges to $k_{0}$, in probability.
+Let $\hat{k}=\arg\max_{1\leq k\leq K}MIC(k)$. Then $\hat{k}$ converges to $k_{0}$, in probability.
 Remark 1.
@@
 It follows from Assumptions 1–3 that $\nabla_{\theta}CGIC(M(\theta))$, evaluated at the true parameter, is a martingale difference sequence.
-By appealing to some variant of the martingale central limit theorem (Hall and Heyde, 2014, Chapter 3), the proofs of Proposition 1 and Theorem 1 for the IID case can be easily extended to the case of stationary ergodic finite-order Markov processes, with MGICE and T-GIC modified accordingly.
-Mimicking AIC and BIC, we set $C(n,k)$ to $\exp\{-2\#(M_{k})/n\}$ and $n^{-\#(M_{k})/n}$, respectively, where #$\#(M_{k})$ is the number of independently adjusted parameters in model $M_{k}$.
-From Theorem 1, the following corollaries follow immediately.
+By appealing to some variant of the martingale central limit theorem (Hall & Heyde 2014, Chapter 3), the proofs of Proposition 1 and Theorem 1 for the IID case can be easily extended to the case of stationary ergodic finite-order Markov processes, with MGICE and MIC modified accordingly.
+Let us consider the following choices of $C(n,k)$: (1) $\exp\{-2\#(M_{k})/n\}$ and (2) $n^{-\#(M_{k})/n}$,
+where $(M_{k})$ is the number of independently adjusted parameters in model $M_{k}$. From Theorem 1, the following corollaries follow immediately.
 Corollary 1.
 Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model.
-If $K=k_{0}$, under Assumptions 1-8, then $T\textbf{-}GIC1=\exp\{-2\#(M_{k})/n\}\times GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
+If $K=k_{0}$, under Assumptions 1-8, then $MIC1=\exp\{-2\#(M_{k})/n\}\times GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
+Note that MIC1 may fail to provide consistent selection when $K>k_{0}$.
 Corollary 2.
 Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model.
-Under Assumptions 1-8, then $T\textbf{-}GIC2=n^{-\#(M_{k})/n}\times GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
-However, similar to AIC, T-GIC1 may fail to provide consistent selection when $K>k_{0}$.
-In the following sections, we will compare the performance of T-GIC1 and T-GIC2 through simulations and applications.
+Under Assumptions 1-8, then $MIC2=n^{-\#(M_{k})/n}\times GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
+In the following sections, we will compare the performance of MIC1 and MIC2 through simulations and applications.
 4 Simulation Study
-In this section, we assess the efficacy of T-GIC1 and T-GIC2 for models with unnormalized PDFs.
-In Section 4.1, we study the consistency of MGICE for $N\times t$ PDF.
-In Sections 4.2 and 4.3, we evaluate the consistency (or otherwise) of T-GIC1 and T-GIC2 and the MGICE for two true models, namely the AR model and the polynomial regression model, each defined on $\mathbb{R}$ and driven by $N\times t$ noise/errors.
+In this section, we assess the efficacy of MIC1 and MIC2 for models with unnormalized PDFs, in comparison with the three bias-corrected criteria GICc, NCIC1, and NCIC2.
+In Section 4.1, we study the consistency of MGICE for Baker PDF.
+In Sections 4.2 and 4.3, we evaluate the consistency (or otherwise) of MIC1 and MIC2 and the MGICE for two true models, namely the AR model and the polynomial regression model, each defined on $\mathbb{R}$ and driven by Baker noise/errors.
 In Section 4.4, we examine a model with a bvariate von Mises PDF on bounded support.
-Upon replications, the empirical distribution of selected orders by T-GIC’s are obtained, as well as the average MGICE and its standard deviation (SD) for parameters in the true model.
-4.1 $N\times t$ PDF
-First, we evaluate the efficacy of the MGICE for unnormalized PDFs. Consider an $N\times t(\alpha,k)$ PDF defined as the product of a normal PDF and a Student t-PDF, given by
+Upon replications, the empirical distribution of selected orders by MIC’s is obtained, as well as the average MGICE and its standard deviation (SD) for parameters in the true model.
+4.1 Baker PDF
+First, we evaluate the efficacy of MGICE for unnormalized PDFs. Consider the Baker PDF (Baker 2022), $N\times t(\alpha,k)$, defined as the product of a normal PDF and a Student t-PDF:
 $f_{N\times t}{(x\mid\alpha,k)}=C(\theta)\frac{\exp\{-\frac{\alpha x^{2}}{2}\}}{(1+x^{2})^{k}},$ (4.1)
-for $-\infty<x<\infty$, where $\alpha>0$, $k>0$, $\theta=(\alpha,k)^{T}$ and $C(\theta)$ is the normalizing constant.
-$N\times t$ is fat-tailed, with the parameters $\alpha$ and $k$ collectively controlling the tail behavior.
+for $\infty<x<\infty$, where $\alpha>0$, $k>0$, $\theta=(\alpha,k)^{T}$ and $C(\theta)$ is the normalizing constant.
+Baker PDF is fat-tailed, with the parameters $\alpha$ and $k$ collectively controlling the tail behavior.
 Specifically, $\alpha$ represents the scale parameter, and $k$ is the power parameter, which is related to the degrees of freedom $\nu$ of a t-PDF via $k=(\nu+1)/2$.
@@
 Instead of the MLE, we employ MGICE for data fitting, which circumvents the need for normalizing constants.
-Considering the affine transformation $X=(Y-\mu)/s$ in Baker (2022), the PDF of $Y$ is $\frac{1}{s}f_{N\times t}(\frac{y-\mu}{s}|\alpha,k)$, with 4 parameters $(\mu,s,\alpha,k)$, where $-\infty<\mu<\infty$ and $s>0$.
+Considering the affine transformation $X=(Y-\mu)/s$ in Baker (2022), the PDF of $Y$ is $\frac{1}{s}f_{N\times t}(\frac{y-\mu}{s}|\alpha,k)$, with 4 parameters $(\mu,s,\alpha,k)$, where $\infty<\mu<\infty$ and $s>0$.
 Let $\{y_{i},i=1,2,\cdots,n\}$ represent $n$ observations and denote $f_{y_{i}}(\theta)=\frac{1}{s}f_{N\times t}(\frac{y_{i}-\mu}{s}|\alpha,k)$.
@@
 Samples are generated using the acceptance-rejection method, by generating random numbers from the $N(0,1/\alpha)$ distribution.
-For numerical optimization in MGICE, we use the Adaptive Moment Estimation (Adam) algorithm (Kingma and Ba, 2015) to jointly optimize all parameters.
-Similar to Baker (2022), for all experiments involving the $N\times t$ PDF, we use the sample mean and standard deviation as starting values for $\mu$ and $s$, and regular initial values for $\alpha$ and $k$, e.g.
+For numerical optimization in MGICE, we use the Adaptive Moment Estimation (Adam) algorithm (Kingma & Ba 2015) to jointly optimize all parameters.
+Similar to Baker (2022), for all experiments involving the Baker PDF, we use the sample mean and standard deviation as starting values for $\mu$ and $s$, and regular initial values for $\alpha$ and $k$, e.g.
 $\alpha=0.25$ or 1, and $k=1$ or 2.
 Table 3 presents the average MGICE and its SD, showing good overall consistency of MGICE, although there is still room for improvement for the parameters $s^{*},\alpha^{*},$ and $k^{*}$.
-Table 4: The average MGICE and its SD for $N\times t$ PDF.
-$N$ Parameter MGICE (SD)
+Table 3: The average MGICE and its SD for Baker PDF.
+NN Parameter MGICE (SD)
 $\mu$ $s$ $\alpha$ $k$
@@
 5000 0.30 (0.01) 0.53 (0.04) 0.48 (0.10) 1.70 (0.14)
-4.2 AR Model with $N\times t$ Noise
-Now, we consider an AR model with $N\times t$ distributed noise and use T-GIC method to select the model order. The stationary mean-centered AR model of order p with $N\times t$ noise is given by
+4.2 AR Model with Baker Noise
+Now, we consider an AR model with Baker distributed noise and use MIC method to select the model order. The stationary mean-centered AR model of order p with Baker noise is given by
 $X_{t}-c=a_{1}(X_{t-1}-c)+\cdots+a_{p}(X_{t-p}-c)+s\varepsilon_{t},$ (4.4)
-where $-\infty<a_{1},\cdots,a_{p},c<\infty$, $s>0$, and $\varepsilon_{t}\sim N\times t(\alpha,k),$ identically and independently.
+where $\infty<a_{1},\cdots,a_{p},c<\infty$, $s>0$, and $\varepsilon_{t}\sim N\times t(\alpha,k),$ identically and independently.
 Let $\{x_{t},t=1,2,\cdots,N\}$ denote $N$ observations from the above model and denote the parameter as $\theta=(a_{1},\cdots,a_{p},c,s,\alpha,k)^{T}$.
@@
 $GIC_{N}(\theta)=\sum_{t=1}^{L}W(x_{t},f_{x_{t}}(\theta))+\sum_{t=L+1}^{N}W(x_{t},f_{x_{t}}(\theta)).$ (4.5)
-where $W(x_{t},f_{x_{t}})=-||\nabla_{x_{t}}\log(f_{x_{t}})||^{2}-2\Delta_{x_{t}}\log(f_{x_{t}})$.
-Discarding the first sum because $W(x_{t},f_{x_{t}}(\theta))$ for $t=1,\cdots L$ are unavailable in the AR(L) model and denoting $n=N-L$, we have
+[…]
+Discarding the first sum because $W(x_{t},f_{x_{t}}(\theta))$ for $t=1,\cdots L$ are unavailable in the AR(L) model and denoting n=NLn=N-L, we have
 $GIC(\theta)$ $=\frac{1}{n}\sum_{t=L+1}^{N}W(x_{t},f_{x_{t}}(\theta)).$ (4.6)
-We conduct 100 replications to obtain the frequency distribution of the selected model orders, ranging from 1 to 10, using T-GIC1 and T-GIC2, with sample size $N=1000,3000,$ and $5000$.
-The true model order $p^{*}$ is set to 3, with true parameters $(a_{1}^{*},a_{2}^{*},a_{3}^{*},c^{*},s^{*},\alpha^{*},k^{*})$ specified at $(0.50,-0.25,0.10,3.00,0.50,0.50,1.50)$.
+We conduct 100 replications to obtain the frequency distribution of the selected model orders, ranging from 1 to 10, using GICc, MIC1 and MIC2, with sample size $N=1000,3000,$ and $5000$.
+The true model order $p^{*}$ is set to 3, with true parameters $(a_{1}^{*},a_{2}^{*},a_{3}^{*},c^{*},s^{*},\alpha^{*},k^{*})$ specified at (0.50,0.25,0.10,3.00,0.50,0.50,1.50)(0.50,-0.25,0.10,3.00,0.50,0.50,1.50).
+Here, NCIC1 and NCIC2 are not considered in this experiment, as they are developed under the assumption of independent and identically distributed observations, and their applicability to dependent data in stationary AR models remains unclear.
 In each replication, the first 200 data are discarded to ensure stationarity.
-Table 5 shows that T-GIC2 tends to underestimate the model order for smaller $N$ but increasingly selects the correct order as $N$ grows.
+Table 4 shows that MIC2 tends to underestimate the model order for smaller $N$ but increasingly selects the correct order as $N$ grows, outperforming the other criteria overall.
 The underfitting for small $N$ is not unexpected due to the small value of $a_{3}^{*}$.
-In contrast, T-GIC1 overestimates the model order with high probability even for large $N$.
+In contrast, MIC1 overestimates the model order with high probability even for large $N$.
 Furthermore, Table 5 reports the average MGICE and its SD for the parameters in the true model, suggesting strong consistency, especially for the AR coefficients and the mean-centering parameter $c$, when $N$ is large.
 Table 4: The frequency distribution of selected orders for the AR(p) model.
-$N$ Method Selected model order $p$
+NN Method Selected model order $p$
 1 2 3 4 5 6 7 8 9 10
-1000 T-GIC1 0 3 22 13 11 1 6 13 14 17
-T-GIC2 1 12 52 9 7 0 4 2 6 7
-3000 T-GIC1 0 0 28 6 7 9 9 5 16 20
-T-GIC2 0 1 81 11 0 4 1 0 0 2
-5000 T-GIC1 0 0 29 10 8 7 9 10 7 20
-T-GIC2 0 0 88 5 2 1 2 1 0 1
-Table 6: The average MGICE and its SD for the AR(3) model with $N\times t$ noise.
-$N$ Parameter MGICE (SD)
+1000 GICc 0 13 57 18 6 1 3 2 0 0
+MIC1 0 3 22 13 11 1 6 13 14 17
+MIC2 1 12 52 9 7 0 4 2 6 7
+3000 GICc 0 0 71 16 2 4 3 2 1 1
+MIC1 0 0 28 6 7 9 9 5 16 20
+MIC2 0 1 81 11 0 4 1 0 0 2
+5000 GICc 0 0 74 16 3 2 2 2 0 1
+MIC1 0 0 29 10 8 7 9 10 7 20
+MIC2 0 0 88 5 2 1 2 1 0 1
+Table 5: The average MGICE and its SD for the AR(3) model with Baker noise.
+NN Parameter MGICE (SD)
 $a_{1}$ $a_{2}$ $a_{3}$ $c$ $s$ $\alpha$ $k$
@@
 5000 0.50 (0.02) -0.25 (0.02) 0.10 (0.02) 3.00 (0.01) 0.52 (0.07) 0.46 (0.10) 1.67 (0.31)
-4.3 Polynomial Regression Model with $N\times t$ Distributed Errors
-Next, we consider a polynomial regression model with $N\times t$ distributed errors and use T-GIC to select the model. The polynomial regression model of degree p is given by
+4.3 Polynomial Regression Model with Baker Distributed Errors
+Next, we consider a polynomial regression model with Baker distributed errors and use MIC to select the model. The polynomial regression model of degree p is given by
 $y=\beta_{1}x+\cdots+\beta_{p}x^{p}+c+s\varepsilon,$ (4.7)
-where $-\infty<\beta_{1},\cdots,\beta_{p},c<\infty$, $s>0$, and $\varepsilon\sim N\times t(\alpha,k),$ identically and independently.
+where $\infty<\beta_{1},\cdots,\beta_{p},c<\infty$, $s>0$, and $\varepsilon\sim N\times t(\alpha,k),$ identically and independently.
 Let $\{(x_{i},y_{i}),i=1,2,\cdots,n\}$ denote $n$ observations from the above model and denote the parameter as $\theta=(\beta_{1},\cdots,\beta_{p},c,s,\alpha,k)^{T}$.
@@
 For each model, from equations (2.1), (2.4), and (2.5), $GIC$ and its MGICE are obtained.
-We conduct 100 replications to obtain the frequency distribution of the selected model degree $p$, ranging from 1 to 10, using T-GIC1 and T-GIC2, with sample size $n=1000,3000,$ and $5000$.
-The true model order $p^{*}$ is set to 3, with true parameters $(\beta_{1}^{*},\beta_{2}^{*},\beta_{3}^{*},c^{*},s^{*},\alpha^{*},k^{*})$ specified as $(-1.5,2.0,5.0,3.0,0.5,0.5,1.5)$.
-Table 7 suggests that both T-GIC1 and T-GIC2 consistently select the correct degree in probability as $n$ increases.
-Moreover, Table 8 reports the average MGICE and its SD for the parameters in the true model, showing good consistency as $n$ increases, although at a slightly slower rate for $\alpha$.
+We conduct 100 replications to obtain the frequency distribution of the selected model degree $p$, ranging from 1 to 10, using GICc, NCIC1, NCIC2, MIC1 and MIC2, with sample size $n=300,500,1000,3000,$ and $5000$.
+The true model order $p^{*}$ is set to 3, with true parameters $(\beta_{1}^{*},\beta_{2}^{*},\beta_{3}^{*},c^{*},s^{*},\alpha^{*},k^{*})$ specified as (1.5,2.0,5.0,3.0,0.5,0.5,1.5)(-1.5,2.0,5.0,3.0,0.5,0.5,1.5).
+For NCIC1 and NCIC2, noise samples of the same size are generated from the classically estimated polynomial regression model distribution using the least squares method under the Gaussian errors assumption.
+Table 6 suggests that both MIC1 and MIC2 consistently select the correct model degree with increasing sample size.
+Although MIC2 shows slightly lower accuracy than NCIC2 when the sample size is small, its performance remains satisfactory and becomes comparable to that of NCIC2 as $n$ increases, eventually achieving the best performance among the criteria.
+Moreover, MIC2 exhibits substantially higher computational efficiency than NCIC2.
+In addition to the theoretical complexity comparison in Section 3.2, we empirically compare the running time ratio of the two methods across different sample sizes, with running time averaged over 10 simulation replications.
+The experiments were conducted using Python 3.10 on a computer with an Intel Xeon Gold 6133 CPU.
+The results, reported in Figure 1, indicate that NCIC2 not only requires more running time than MIC2 but also shows a much faster growth in running time as the sample size increases.
+Table 7 reports the average MGICE and its SD for the parameters in the true model, showing good consistency as $n$ increases, although at a slightly slower rate for $\alpha$.
+Figure 1: Empirical running time ratios between NCIC2 and MIC2 for the polynomial regression model across different sample sizes.
 Table 6: The frequency distribution of selected degrees for the polynomial regression model.
@@
 1 2 3 4 5 6 7 8 9 10
-1000 T-GIC1 0 0 94 6 0 0 0 0 0 0
-T-GIC2 0 0 97 3 0 0 0 0 0 0
-3000 T-GIC1 0 0 96 4 0 0 0 0 0 0
-T-GIC2 0 0 96 4 0 0 0 0 0 0
-5000 T-GIC1 0 0 97 3 0 0 0 0 0 0
-T-GIC2 0 0 97 3 0 0 0 0 0 0
+300 GICc 0 0 67 23 4 2 3 0 1 0
+NCIC1 8 0 68 10 7 2 2 1 1 1
+NCIC2 0 0 96 4 0 0 0 0 0 0
+MIC1 0 0 77 14 6 1 2 0 0 0
+MIC2 0 0 81 11 5 1 2 0 0 0
+500 GICc 0 0 85 9 4 1 1 0 0 0
+NCIC1 4 1 66 15 8 1 0 1 3 1
+NCIC2 0 0 91 9 0 0 0 0 0 0
+MIC1 0 0 84 9 5 1 1 0 0 0
+MIC2 0 0 87 7 4 1 1 0 0 0
+1000 GICc 0 0 93 7 0 0 0 0 0 0
+NCIC1 1 0 78 11 3 2 2 0 2 1
+NCIC2 0 0 96 4 0 0 0 0 0 0
+MIC1 0 0 94 6 0 0 0 0 0 0
+MIC2 0 0 97 3 0 0 0 0 0 0
+3000 GICc 0 0 97 3 0 0 0 0 0 0
+NCIC1 0 0 77 12 5 2 2 2 0 0
+NCIC2 0 0 98 2 0 0 0 0 0 0
+MIC1 0 0 96 4 0 0 0 0 0 0
+MIC2 0 0 96 4 0 0 0 0 0 0
+5000 GICc 0 0 97 3 0 0 0 0 0 0
+NCIC1 0 0 86 12 0 1 0 1 0 0
+NCIC2 0 0 91 9 0 0 0 0 0 0
+MIC1 0 0 97 3 0 0 0 0 0 0
+MIC2 0 0 97 3 0 0 0 0 0 0
 Table 7: The average MGICE and its SD for the cubic polynomial regression model.
@@
 $\beta_{1}$ $\beta_{2}$ $\beta_{3}$ $c$ $s$ $\alpha$ $k$
+300 -1.52 (0.47) 2.03 (0.52) 4.95 (0.46) 2.95 (0.53) 0.90 (0.65) 0.74 (2.66) 3.39 (4.12)
+500 -1.53 (0.34) 2.01 (0.42) 4.94 (0.37) 2.98 (0.35) 0.73 (0.49) 0.25 (0.39) 2.76 (2.56)
 1000 -1.50 (0.03) 2.00 (0.01) 5.00 (0.01) 3.00 (0.02) 0.56 (0.21) 0.33 (0.20) 2.08 (1.27)
@@
 4.4 A Bivariate Model with a Von Mises PDF
-As a final case, we use T-GIC to select the dimension of the parameter space of a model with unnormalized PDF on bounded support.
-Consider the bivariate von Mises PDF (Singh et al., 2002) of two circular random variables $X_{1},X_{2}$, given by
-[…]
+As a final case, we use MIC to select the dimension of the parameter space of a model with unnormalized PDF on bounded support.
+Consider the bivariate von Mises PDF (Singh et al.
+2002) of two circular random variables $X_{1},X_{2}$, given by
+$f(x_{1},x_{2}\mid\theta)=C(\theta)\exp\{$ $\kappa_{1}\cos(x_{1}-\mu_{1})+\kappa_{2}\cos(x_{2}-\mu_{2})$
 $+\lambda\sin(x_{1}-\mu_{1})\sin(x_{2}-\mu_{2})\},$ (4.8)
-for $0\leq x_{1}$, $x_{2}<2\pi$, where $\kappa_{1}$, $\kappa_{2}\geq 0$, $0\leq\mu_{1}$, $\mu_{2}<2\pi$, $-\infty<\lambda<\infty$, $\theta=(\kappa_{1},\kappa_{2},\mu_{1},\mu_{2},\lambda)^{T}$ and $C(\theta)$ is the normalizing constant.
+for $0\leq x_{1}$, $x_{2}<2\pi$, where $\kappa_{1}$, $\kappa_{2}\geq 0$, $0\leq\mu_{1}$, $\mu_{2}<2\pi$, $\infty<\lambda<\infty$, $\theta=(\kappa_{1},\kappa_{2},\mu_{1},\mu_{2},\lambda)^{T}$ and $C(\theta)$ is the normalizing constant.
 The parameter $\lambda$ quantifies the dependency between two circular random variables $X_{1},X_{2}$.
@@
 For each model, from equations (2.1), (2.4), and (2.5), $GIC$ and its MGICE are obtained.
-We conduct 100 replications to evaluate the performance of model selection between $m_{1}$ and $m_{2}$, using T-GIC1 and T-GIC2, with sample size $n=300,500,\text{and }1000$.
+We conduct 100 replications to evaluate the performance of model selection between $m_{1}$ and $m_{2}$, using GICc, NCIC1, NCIC2, MIC1 and MIC2, with sample size $n=300,500,\text{and }1000$.
 The true parameter dimension is set to 5, with true parameter $\theta^{*}=(2.0,1.0,1.5,2.5,3.0)^{T}$.
-Samples are generated using the acceptance-rejection method, by generating a random number from the uniform distribution on $[0,2\pi)\times[0,2\pi)$.
-For numerical optimization in MGICE, we use the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm (Nocedal and Wright, 2006) to jointly optimize all parameters, with all parameters initialized to zero.
-In all replications, both T-GIC1 and T-GIC2 correctly select the true parameter dimension 5.
+Samples are generated using the acceptance-rejection method, by generating a random number from the uniform distribution on $0,2\pi)\times[0,2\pi)$.
+For NCIC1 and NCIC2, noise samples of the same size are generated from the same uniform distribution.
+For numerical optimization in MGICE, we use the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm (Nocedal & Wright 2006) to jointly optimize all parameters, with all parameters initialized to zero.
+In all replications, all criteria correctly select the true parameter dimension 5.
+However, NCIC2 requires substantially more running time than MIC2, over ten times longer when the sample size exceeds 3000, as shown in Figure  2.
 Table 8 presents the average MGICE and its SD for parameters in the true model $m_{2}$, demonstrating the strong consistency of MGICE especially when $n$ is large.
+Figure 2: Empirical running time ratios between NCIC2 and MIC2 for the bivariate model with a Von Mises PDF across different sample sizes.
 Table 8: The average MGICE and its SD for the model $m_{2}$.
@@
 5 Real Data
-In this section, we apply T-GIC for model selection with real data across three domains.
-Now, Baker (2022) has argued that the $N\times t$ PDF provides a more realistic framework for modeling real-world data, especially in finance, and serves as an effective tool for assessing robustness and performing sensitivity analyses.
-In Sections 5.1 and 5.2, we consider fitting autoregression to some finance data, and polynomial regression to some car data, with $N\times t$ noise/errors, using T-GIC1 and T-GIC2 for model selection.
+In this section, we apply MIC for model selection with real data across three domains.
+Now, Baker (2022) has argued that the Baker PDF provides a more realistic framework for modeling real-world data, especially in finance, and serves as an effective tool for assessing robustness and performing sensitivity analyses.
+In Sections 5.1 and 5.2, we consider fitting autoregression to some finance data, and polynomial regression to some car data, with Baker noise/errors, using MIC1 and MIC2 for model selection.
 We also compare the results with those based on AIC and BIC with Gaussian noise/errors.
@@
 5.1 Finance Data
-We analyze the logged returns on the Financial Times Stock Exchange 100 Index (FTSE 100), covering 9013 observations from January 1986 to April 2021. Figure 1 shows its time plot.
-Figure 1: Time plot of logged returns on the FTSE 100 from January 1986 to April 2021.
-Baker (2022) assumed, for simplicity, independent data and fitted an $N\times t$ distribution.
+We analyze the logged returns of three stock market indices: the Financial Times Stock Exchange 100 Index (FTSE 100) with 9013 observations from January 1986 to April 2021, the Nikkei Stock Average (Nikkei 225) with 5159 observations from July 2003 to August 2024, and the Standard & Poor’s 500 Index (S&P 500) with 3827 observations from January 2010 to March 2025.
+Their time plots are displayed in Figure 3.
+(a) FTSE 100
+(b) Nikkei 225
+(c) S&P 500
+Figure 3: Time plots of logged returns on the FTSE 100, Nikkei 225, and S&P 500.
+Baker (2022) assumed, for simplicity, independent data and fitted a Baker distribution.
 Instead, we consider an AR model to allow for data dependence.
-For comparison, we use T-GIC1 and T-GIC2 to select an AR model with $N\times t$ noise in equation (4.4), and AIC and BIC to select an AR model with Gaussian noise, given by
+For comparison, we use MIC1 and MIC2 to select an AR model with Baker noise in equation (4.4), and AIC and BIC to select an AR model with Gaussian noise, given by
 $X_{t}-c=a_{1}(X_{t-1}-c)+\cdots+a_{p}(X_{t-p}-c)+\varepsilon_{t},$ (5.1)
 where $\varepsilon_{t}\sim N(0,\sigma^{2})$.
-Tables 10 and 11 summarize the results.
-Under Gaussian noise, AIC selects a seventh-order model, while BIC selects a sixth-order model.
-However, under $N\times t$ noise, T-GIC1 selects a sixth-order AR model, while T-GIC2 selects a second-order AR model, both yielding estimates of $\alpha$ and $k$ similar to Baker (2022).
-Since AR models offer a convenient framework for prediction, Table 12 presents a comparison of out-of-sample performance between the two fitted models, either chosen by AIC vs BIC or by T-GIC1 vs T-GIC2.
-In each case, we use the first part of the dataset for fitting and the remaining 100 data for the rolling $m$-step-ahead forecast.
-It is interesting to observe, for T-GIC1 vs T-GIC2, that the fitted AR(6) model performs better in forecasting for the short term than the fitted AR(2) model but worse for the longer term.
-However, for AIC vs BIC, the fitted AR(6) performs better than the fitted AR(7) uniformly.
-Overall, Table 12 suggests that the effect of heavy-tailed innovation kicks in for the longer term forecasting.
+Tables 9 and 10 summarize the results.
+Under Gaussian noise, AIC selects a seventh-order model, while BIC selects a sixth-order model for the FTSE 100.
+For the Nikkei 225, both AIC and BIC select a first-order model, while for the S&P 500, both select a ninth-order model.
+However, under Baker noise, MIC1 selects a sixth-order AR model, while MIC2 selects a second-order AR model for the FTSE 100, yielding estimates of $\alpha$ and $k$ similar to Baker (2022).
+For the Nikkei 225, MIC1 selects a sixth-order AR model, whereas MIC2 favors a first-order AR model.
+For the S&P 500, both MIC1 and MIC2 select a first-order AR model.
 Table 9: Selection results for AR model with two noise PDFs.
 $\varepsilon_{t}\sim N(0,\sigma^{2})$
-$p$ 1 2 3 4 5 6 7 8 9 10
-AIC -55538.20 -55543.89 -55557.30 -55571.54 -55574.02 -55586.90 -55593.47 -55592.42 -55591.97 -55590.44
+pp 1 2 3 4 5 6 7 8 9 10
+FTSE 100 AIC -55538.20 -55543.89 -55557.30 -55571.54 -55574.02 -55586.90 -55593.47 -55592.42 -55591.97 -55590.44
 BIC -55516.88 -55515.47 -55521.77 -55528.91 -55524.27 -55530.05 -55529.51 -55521.36 -55513.80 -55505.16
+Nikkei 225 AIC -29127.55 -29125.80 -29127.32 -29126.13 -29126.43 -29124.65 -29122.66 -29120.81 -29118.96 -29118.56
+BIC -29107.90 -29099.60 -29094.58 -29086.84 -29080.59 -29072.27 -29063.72 -29055.32 -29046.92 -29039.98
+S&P 500 AIC -23789.81 -23801.46 -23800.63 -23808.10 -23806.12 -23825.54 -23854.07 -23867.60 -23879.13 -23877.23
+BIC -23771.06 -23776.47 -23769.38 -23770.60 -23762.38 -23775.54 -23797.82 -23805.10 -23810.38 -23802.23
 $\varepsilon_{t}\sim N\times t(\alpha,k)$
-$p$ 1 2 3 4 5 6 7 8 9 10
-T-GIC1 12716.15 12837.14 12828.04 12812.87 12821.58 12866.07 12787.31 12862.53 12746.68 12754.23
-T-GIC2 12706.13 12816.92 12797.73 12772.53 12771.13 12805.35 12716.93 12781.66 12656.55 12654.07
+pp 1 2 3 4 5 6 7 8 9 10
+FTSE 100 MIC1 12716.15 12837.14 12828.04 12812.87 12821.58 12866.07 12787.31 12862.53 12746.68 12754.23
+MIC2 12706.13 12816.92 12797.73 12772.53 12771.13 12805.35 12716.93 12781.66 12656.55 12654.07
+Nikkei 225 MIC1 7589.51 7593.72 7602.29 7590.82 7595.97 7606.32 7545.75 7568.10 7587.28 7602.21
+MIC2 7579.88 7574.47 7573.40 7552.38 7547.92 7548.61 7479.00 7491.63 7501.09 7506.32
+S&P 500 MIC1 2404.78 2348.69 2395.95 2132.21 2189.68 2339.97 2330.62 2305.78 2198.37 2381.86
+MIC2 2400.86 2341.04 2384.24 2118.33 2171.88 2317.16 2304.13 2275.85 2166.30 2343.28
 Table 10: Parameter estimate results for the selected AR model.
@@
 Order (method) $(a_{1},\cdots,a_{p})$ $c$ $\sigma^{2}$
-$p=7$ (AIC) (-0.0004, -0.0266, -0.0448, 0.0426, -0.0215, -0.0406, 0.0308) 0.0002 0.0001
+FTSE 100 $p=7$ (AIC) (-0.0004, -0.0266, -0.0448, 0.0426, -0.0215, -0.0406, 0.0308) 0.0002 0.0001
 $p=6$ (BIC) (-0.0016, -0.0272, -0.0435, 0.0412, -0.0223, -0.0406) 0.0002 0.0001
+Nikkei 225 $p=1$ (AIC, BIC) (-0.0309) 0.0003 0.0002
+S&P 500 $p=9$ (AIC, BIC) (-0.0926, 0.0487, -0.0177, -0.0449, -0.0143, -0.0601, 0.0795, -0.0580, 0.0594) 0.0004 0.0001
 $\varepsilon_{t}\sim N\times t(\alpha,k)$
 Order (method) $(a_{1},\cdots,a_{p})$ $c$ s $\alpha$ $k$
-$p=6$ (T-GIC1) (-0.0212, -0.0813, -0.0158, -0.0265, -0.0230, -0.0331) 0.0006 0.0129 0.0104 2.0799
-$p=2$ (T-GIC2) (-0.0256, -0.0877) 0.0007 0.0129 0.0104 2.0807
+FTSE 100 $p=6$ (MIC1) (-0.0212, -0.0813, -0.0158, -0.0265, -0.0230, -0.0331) 0.0006 0.0129 0.0104 2.0799
+$p=2$ (MIC2) (-0.0256, -0.0877) 0.0007 0.0129 0.0104 2.0807
+Nikkei 225 $p=6$ (MIC1) (-0.0445, -0.0020, -0.0260, -0.0157, -0.0303, 0.0031) 0.0008 0.0115 0.1030 1.2176
+$p=1$ (MIC2) (-0.0479) 0.0011 0.0116 0.1052 1.2292
+S&P 500 $p=1$ (MIC1, MIC2) (-0.0168) 0.0007 0.0525 0.0093 1.9206
+Since AR models offer a convenient framework for prediction, Table 11 presents a comparison of out-of-sample performance between the fitted models, either chosen by AIC vs BIC or by MIC1 vs MIC2.
+In each case, we use the first part of the dataset for fitting and the remaining 100 data for the rolling $m$-step-ahead forecast.
+It is interesting to observe that, for the FTSE 100, the fitted AR(6) model selected by MIC1 performs better in forecasting for the short term than the fitted AR(2) model selected by MIC2 but worse for the longer term.
+However, for AIC vs BIC, the fitted AR(6) performs better than the fitted AR(7) uniformly.
+The results suggest that the effect of heavy-tailed innovation kicks in for the longer term forecasting.
+For the Nikkei 225, the fitted AR(6) model chosen by MIC2 achieves the best forecasting performance.
+For the S&P 500, the fitted AR(1) model selected by both MIC1 and MIC2, which involves heavy-tailed innovation, uniformly outperforms the fitted AR(9) model selected by AIC and BIC.
 Table 11: MSE and the ratio in the rolling m-step-ahead forecast on AR models. (Row-wise minimum values are underlined.)
-$m$ $\varepsilon_{t}\sim N\times t(\alpha,k)$ $\varepsilon_{t}\sim N(0,\sigma^{2})$
+Data $m$ $\varepsilon_{t}\sim N\times t(\alpha,k)$ $\varepsilon_{t}\sim N(0,\sigma^{2})$
 $MSE(AR(2))$ $MSE(AR(6))$ $\frac{MSE(AR(2))}{MSE(AR(6))}$ $MSE(AR(6))$ $MSE(AR(7))$ $\frac{MSE(AR(6))}{MSE(AR(7))}$
-1 $8.2026\times 10^{-5}$ $8.1961\times 10^{-5}$ 1.0008 $8.1247\times 10^{-5}$ $8.1534\times 10^{-5}$ 0.9965
+FTSE 100 1 $8.2026\times 10^{-5}$ $8.1961\times 10^{-5}$ 1.0008 $8.1247\times 10^{-5}$ $8.1534\times 10^{-5}$ 0.9965
 2 $8.2508\times 10^{-5}$ $8.2340\times 10^{-5}$ 1.0020 $8.1492\times 10^{-5}$ $8.1763\times 10^{-5}$ 0.9967
@@
 5 $8.2529\times 10^{-5}$ $8.3208\times 10^{-5}$ 0.9918 $8.4179\times 10^{-5}$ $8.4545\times 10^{-5}$ 0.9957
+$MSE(AR(1))$ $MSE(AR(6))$ $\frac{MSE(AR(1))}{MSE(AR(6))}$ $MSE(AR(1))$ - -
+Nikkei 225 1 $4.3790\times 10^{-4}$ $4.3024\times 10^{-4}$ 1.0178 $4.3970\times 10^{-4}$ - -
+2 $4348\times 10^{-4} 4.4$$4128\times 10^{-4}$ 1.0050 $4.4373\times 10^{-4}$ - -
+3 $4705\times 10^{-4} 4.4$$4556\times 10^{-4}$ 1.0033 $4.4671\times 10^{-4}$ - -
+4 $5225\times 10^{-4} 4.4$$4758\times 10^{-4}$ 1.0104 $4.5128\times 10^{-4}$ - -
+5 $4.5423\times 10^{-4}$ $4.5263\times 10^{-4}$ 1.0035 $4.5522\times 10^{-4}$ - -
+$MSE(AR(1))$ - - $MSE(AR(9))$ - -
+S&P 500 1 $8073\times 10^{-5} - - 8.8$$8697\times 10^{-5} - -$
+2 $8.9475\times 10^{-5}$ - - $8.9767\times 10^{-5}$ - -
+3 $9.0349\times 10^{-5}$ - - $9.1379\times 10^{-5}$ - -
+4 $9.1280\times 10^{-5}$ - - $9.2134\times 10^{-5}$ - -
+5 $9.2046\times 10^{-5}$ - - $9.4757\times 10^{-5}$ - -
 5.2 Car Data
@@
 The data plot suggests a nonlinear relationship between mpg and horsepower.
-Gareth et al.
+James et al.
 (2013) fitted a polynomial regression model with Gaussian errors to this dataset.
-In this study, we apply a polynomial regression model with $N\times t$ errors and use the T-GIC1 and T-GIC2 to select the degree of the polynomial, setting candidate degrees from 1 to 10.
+In this study, we apply a polynomial regression model with Baker errors and use the MIC1 and MIC2 to select the degree of the polynomial, setting candidate degrees from 1 to 10.
 For comparison, we also use AIC and BIC with Gaussian errors.
-For $N\times t$ errors we employ the model in equation (4.7), and for Gaussian errors we employ the model given by
+For Baker errors we employ the model in equation (4.7), and for Gaussian errors we employ the model given by
 $y=\beta_{1}x+\cdots+\beta_{p}x^{p}+c+\varepsilon,$ (5.2)
 where $\varepsilon\sim N(0,\sigma^{2})$.
-Tables 13 and 14 report model selection and parameter estimation, respectively.
+Tables 12 and 13 report model selection and parameter estimation, respectively.
 Under Gaussian errors, AIC selects a seventh-degree polynomial regression model, while BIC selects a quadratic polynomial regression model.
-In contrast, both T-GIC1 and T-GIC2 select a quadratic polynomial regression model with $N\times t$ errors, and give estimates similar to those based on the BIC with Gaussian errors.
+In contrast, both MIC1 and MIC2 select a quadratic polynomial regression model with Baker errors, and give estimates similar to those based on the BIC with Gaussian errors.
 This suggests that a quadratic model chosen by BIC is an appropriate model rather than the seventh-degree model chosen by AIC.
-This example shows that T-GIC1 and T-GIC2 can be used to check the choice by AIC and BIC, even when the latter can be applied, and highlights an additional utility of T-GIC coupled with the $N\times t$ PDF.
+This example shows that MIC1 and MIC2 can be used to check the choice by AIC and BIC, even when the latter can be applied, and highlights an additional utility of MIC coupled with the Baker PDF.
 Table 12: Selection results for polynomial regression model with two error PDFs.
 $\varepsilon\sim N(0,\sigma^{2})$
-$p$ 1 2 3 4 5 6 7 8 9 10
+pp 1 2 3 4 5 6 7 8 9 10
 AIC -186.29 -242.96 -243.44 -242.01 -249.08 -248.42 -250.44 -248.80 -248.39 -247.24
@@
 $\varepsilon\sim N\times t(\alpha,k)$
-$p$ 1 2 3 4 5 6 7 8 9 10
-T-GIC1 27.69 34.24 34.07 33.89 33.08 29.71 32.66 28.84 32.38 31.93
-T-GIC2 27.41 33.55 33.05 32.55 31.45 27.95 30.42 26.60 29.56 28.86
+pp 1 2 3 4 5 6 7 8 9 10
+MIC1 27.69 34.24 34.07 33.89 33.08 29.71 32.66 28.84 32.38 31.93
+MIC2 27.41 33.55 33.05 32.55 31.45 27.95 30.42 26.60 29.56 28.86
 Table 13: Parameter estimate results for the selected polynomial regression model.
@@
 Degree (method) $(\beta_{1},\cdots,\beta_{p})$ $c$ $\sigma^{2}$
-$p=7$ (AIC) (-0.2966, 0.0723, -0.1232, 0.0412, 0.0489, -0.0321, 0.0051) 3.0376 0.0296
+$p=7$ (AIC)
+(-0.2966, 0.0723, -0.1232, 0.0412, 0.0489, -0.0321, 0.0051)
+3.0376 0.0296
 $p=2$ (BIC) (-0.3448,0.0578) 3.0407 0.0309
@@
 Degree (method) $(\beta_{1},\cdots,\beta_{p})$ $c$ $s$ $\alpha$ $k$
-$p=2$ (T-GIC1, T-GIC2) (-0.3838, 0.0800) 3.0288 0.3757 0.4973 3.2389
+$p=2$ (MIC1, MIC2) (-0.3838, 0.0800) 3.0288 0.3757 0.4973 3.2389
 We also examine the skewness and excess kurtosis of the residuals and consider the matching between the theoretical values and their sample estimates based on the fitted residuals, as shown in Table 14.
@@
 For the fitted models with Gaussian errors, the skewness and the excess kurtosis of the errors are both theoretically 0.
-On the other hand, for $N\times t(\alpha,k)$ errors, the skewness is theoretically 0, but the excess kurtosis would not be 0 unless $k=0$.
-Since it is difficult to calculate the excess kurtosis for $N\times t(\alpha,k)$ PDF due to its intractable normalizing constant, we sample from $N\times t$ PDF and use the mean of the sample estimates from 1000 replications as the theoretical value, along with its SE.
-Apparently, the fitted residuals of the models chosen by AIC, BIC, T-GIC1 and T-GIC2 have all produced very small negative skewness of similar size, matching their theoretical value of zero reasonably well.
-However, for excess kurtosis, while the matching is far from being satisfactory for the Gaussian models chosen by AIC and BIC, the matching for models chosen by T-GIC1 and T-GIC2 coupled with $N\times t$ PDFs is much better.
+On the other hand, for Baker errors, the skewness is theoretically 0, but the excess kurtosis would not be 0 unless $k=0$.
+Since it is difficult to calculate the excess kurtosis for Baker PDF due to its intractable normalizing constant, we sample from Baker PDF and use the mean of the sample estimates from 1000 replications as the theoretical value, along with its SE.
+Apparently, the fitted residuals of the models chosen by AIC, BIC, MIC1 and MIC2 have all produced very small negative skewness of similar size, matching their theoretical value of zero reasonably well.
+However, for excess kurtosis, while the matching is far from being satisfactory for the Gaussian models chosen by AIC and BIC, the excess kurtosis from the fitted model chosen by MIC1 and MIC2 coupled with Baker PDF is 1.29 (SE: 0.21) versus its residual counterpart of 1.33 (SE: 0.54), which is much better.
 Table 14: Comparison of residual skewness and excess kurtosis.
-Noise Model Skewness Excess kurtosis
+Noise Degree Skewness Excess kurtosis
 Theoretical value Sample estimate (SE) Theoretical value Sample estimate (SE)
-$N(0,\sigma^{2})$ AR(7) 0 -0.08 (0.22) 0 1.13 (0.59)
-AR(2) 0 -0.14 (0.20) 0 0.87 (0.49)
-$N\times t(\alpha,k)$ AR(2) 0 -0.22 (0.23) 1.29 (0.21) 1.33 (0.54)
+$N(0,\sigma^{2})$ $p=7$ 0 -0.08 (0.22) 0 1.13 (0.59)
+$p=2$ 0 -0.14 (0.20) 0 0.87 (0.49)
+$N\times t(\alpha,k)$ $p=2$ 0 -0.22 (0.23) 1.29 (0.21) 1.33 (0.54)
 5.3 Wind Direction Data
-Finally, we fit a bivariate model with von Mises PDF, as described in section 4.3, to some wind direction data, using T-GIC1 and T-GIC2 for model selection.
+Finally, we fit a bivariate model with von Mises PDF, as described in section 4.3, to some wind direction data, using MIC1 and MIC2 for model selection.
 Here, the wind direction is represented as a circular variable in radians.
@@
 We fit the data using two candidate models, $m_{1}$ and $m_{2}$, representing two scenarios: $\lambda=0$ and $\lambda\in R$ in equation (4.8), respectively.
-For comparison, results for T-GIC1, T-GIC2 are summarized in Table 16.
-For $m_{2}$, both T-GIC1 and T-GIC2 are higher, indicating a better fit than $m_{1}$.
+For comparison, results for MIC1, MIC2 are summarized in Table 15.
+For $m_{2}$, both MIC1 and MIC2 are higher, indicating a better fit than $m_{1}$.
 This suggests that the wind directions at Tokyo on 00:00 and 12:00 in 2023 are dependent, consistent with the results in 2008 by Matsuda et al.
 (2021).
-Table 16: Parameter estimate and T-GIC results for models $m_{1}$ and $m_{2}$.
-Model Order $\kappa_{1}$ $\kappa_{2}$ $\mu_{1}$ $\mu_{2}$ $\lambda$ T-GIC1 T-GIC2
+Table 15: Parameter estimate and MIC results for models $m_{1}$ and $m_{2}$.
+Model Order $\kappa_{1}$ $\kappa_{2}$ $\mu_{1}$ $\mu_{2}$ $\lambda$ MIC1 MIC2
 $m_{1}$ 4 0.4607 0.3057 5.6711 2.3854 0 0.1014 0.0971
@@
 6 Conclusion
-In this paper, based on GIC (Cheng and Tong, 2024), we propose a consistent selection criterion, T-GIC, for nested models with possibly unnormalized PDFs.
-Compared with existing model selection methods, T-GIC offers three advantages.
+In this paper, we propose a consistent selection criterion, MIC, for nested models with possibly unnormalized PDFs.
+Compared with existing model selection methods, MIC offers three advantages.
 First, it can handle both normalized and unnormalized PDFs.
@@
 Third, it reduces significantly computational costs by avoiding the calculation of the normalizing constant and the bias correction.
-To showcase the efficacy of T-GIC for unnormalized PDFs, we have studied AR models and polynomial regression models with $N\times t$ noise/errors, the normalizing constants of which are typically intractable.
-Extensive simulation studies and real data applications have demonstrated consistency and effectiveness of T-GIC.
-Furthermore, we have shown excellent performance of T-GIC on PDFs with bounded support through experiments on models with bivariate von Mises PDFs and real wind direction data.
-We have discussed how to use T-GIC for consistent model selection with PDFs supported on $\mathbb{R}^{d}.$ The simulation results and the real example with wind data strongly suggest that a promising direction for future research is to extend T-GIC to cover PDFs supported on bounded intervals $(a,b)$.
-Another direction of extension is to discrete data, thereby availing T-GIC of the opportunity of selecting an appropriate model, such as an Ising model (Friel, 2013; Everitt et al., 2017), in the area of discrete Markov random fields and spatial statistics.
+To showcase the efficacy of MIC for unnormalized PDFs, we have studied AR models and polynomial regression models with Baker noise/errors, the normalizing constants of which are typically intractable.
+Extensive simulation studies and real data applications have demonstrated consistency and effectiveness of MIC.
+Furthermore, we have shown excellent performance of MIC on PDFs with bounded support through experiments on models with bivariate von Mises PDFs and real wind direction data.
+We have discussed how to use MIC for consistent model selection with PDFs supported on $\mathbb{R}^{d}.$ The simulation results and the real example with wind data strongly suggest that a promising direction for future research is to extend MIC to cover PDFs supported on bounded intervals $(a,b)$.
+Another direction of extension is to discrete data, thereby availing MIC of the opportunity of selecting an appropriate model, such as an Ising model (Friel 2013, Everitt et al.
+2017), in the area of discrete Markov random fields and spatial statistics.
 It is also intriguing to explore the use of combinations of different multiplying factors $C(n,k)$.
-Drawing inspiration from AIC and BIC, we have set $C(n,k)$ to $\exp\{-2\#(M_{k})/n\}$ and $n^{-\#(M_{k})/n}$, respectively.
-Although $\exp\{-2\#(M_{k})/n\}$ may not yield a consistent estimate of the true order, there is significant scope for combining it with $n^{-\#(M_{k})/n}$, similar to existing approaches for combining AIC and BIC in model selection (Hsu et al., 2019).
-Last, but not least, an exploration of T-GIC into non-nested models should be exciting.
-Appendix A Appendix
-A.1 Proof of Proposition 1
+We have set $C(n,k)$ to $\exp\{-2\#(M_{k})/n\}$ and $n^{-\#(M_{k})/n}$, respectively.
+Although $\exp\{-2\#(M_{k})/n\}$ may not yield a consistent estimate of the true order, there is significant scope for combining it with $n^{-\#(M_{k})/n}$, similar to existing approaches developed in Ing (2007).
+Last, but not least, an exploration of MIC into non-nested models should be exciting.
+7 Disclosure Statement
+We declare that we do not have any relevant financial or non-financial competing interests.
+[…]
+[…]
+Proof.
 To simplify notations, we define an average operator, $P_{n}$, for data sample $x_{1},\cdots,x_{n}$ applied to any function $g(x,\cdot)$ by
@@
 Note that $\hat{\alpha}_{0}$ satisfies the equation:
-00 0 $=P_{n}\nabla_{\alpha}W(x,p_{M(\hat{\alpha}_{0},\beta^{*})})$
-$=P_{n}\nabla_{\alpha}W(x,p_{M(\alpha^{*},\beta^{*})})+P_{n}\nabla_{\alpha}\nabla^{T}_{\alpha}W(x,p_{M(\alpha^{*},\beta^{*})})(\hat{\alpha}_{0}-\alpha^{*})+O_{p}(n^{-1})$ (A.1)
+$0$ $=P_{n}\nabla_{\alpha}W(x,p_{M(\hat{\alpha}_{0},\beta^{*})})$
+[…]
 where the $O_{p}(n^{-1})$ term follows from mean value theorem and Assumption 8, and that the MGICE is root-n consistent. Similarly, the unconstrained MGICE satisfies the following equation.
 $0=P_{n}\nabla_{\alpha}W(x,p_{M(\alpha^{*},\beta^{*})})+P_{n}\nabla_{\alpha}\nabla^{T}_{\alpha}W(x,p_{M(\alpha^{*},\beta^{*})})(\hat{\alpha}-\alpha^{*})$
-$+P_{n}\nabla_{\alpha}\nabla^{T}_{\beta}W(x,p_{M(\alpha^{*},\beta^{*})})(\hat{\beta}-\beta^{*})+O_{p}(n^{-1}).$ (A.2)
+[…]
 The preceding two equations imply that
-$\hat{\alpha}_{0}-\alpha^{*}=\hat{\alpha}-\alpha^{*}+D^{-1}_{\alpha^{*},\alpha^{*}}D_{\alpha^{*},\beta^{*}}(\hat{\beta}-\beta^{*})+O_{p}(n^{-1}),$ (A.3)
+[…]
 where $D_{a^{*},b^{*}}=P_{n}\nabla_{a}\nabla^{T}_{b}W(x,p_{M(\alpha^{*},\beta^{*})})$, with $a$, $b$ being either $\alpha$ or $\beta$, and $D_{\alpha^{*},\alpha^{*}}$ is assumed to be invertible.
@@
 $+O_{p}(n^{-3/2})$
-$=\begin{pmatrix}(\hat{\alpha}-\alpha^{*})^{T}&(\hat{\beta}-\beta^{*})^{T}\end{pmatrix}\begin{pmatrix}D_{\alpha^{*},\alpha^{*}}&D_{\alpha^{*},\beta^{*}}\\ D_{\beta^{*},\alpha^{*}}&D_{\beta^{*},\alpha^{*}}D^{-1}_{\alpha^{*},\alpha^{*}}D_{\alpha^{*},\beta^{*}}\end{pmatrix}\binom{\hat{\alpha}-\alpha^{*}}{\hat{\beta}-\beta^{*}}+O_{p}(n^{-3/2})$ (A.4)
+[…]
 where $I$ denotes the identity matrix of dimension $k_{0}$. Similarly, we have
 $[\log(GIC(M(\alpha^{*},\beta^{*})))-\log P_{n}\{W(x,p_{M(\hat{\alpha},\hat{\beta})})\}]\times[2P_{n}\{W(x,p_{M(\alpha^{*},\beta^{*})})\}]$
-$=\begin{pmatrix}(\hat{\alpha}-\alpha^{*})^{T}&(\hat{\beta}-\beta^{*})^{T}\end{pmatrix}\begin{pmatrix}D_{\alpha^{*},\alpha^{*}}&D_{\alpha^{*},\beta^{*}}\\ D_{\beta^{*},\alpha^{*}}&D_{\beta^{*},\beta^{*}}\end{pmatrix}\binom{\hat{\alpha}-\alpha^{*}}{\hat{\beta}-\beta^{*}}+O_{p}(n^{-3/2})$ (A.5)
+[…]
 Note that the matrix in the middle of the quadratic form depends on $n$ and it converges in probability to $D(\theta^{*})$ defined in (3.3) which is a negative-definite matrix.
 We can similarly partition it into a 2 by 2 block matrix:
-$D(\theta^{*})=\begin{pmatrix}D(\alpha^{*},\alpha^{*})&D(\alpha^{*},\beta^{*})\\ D(\beta^{*},\alpha^{*})&D(\beta^{*},\beta^{*})\end{pmatrix}$ (A.6)
-Subtracting (A.5) from (A.4) yields:
+[…]
+[…]
 $[\log P_{n}\{W(x,p_{M(\hat{\alpha},\hat{\beta})})\}-\log P_{n}\{W(x,p_{M(\hat{\alpha}_{0},\beta^{*})})\}]\times[2P_{n}\{W(x,p_{M(\alpha^{*},\beta^{*})})\}]$
-$=-(\hat{\beta}-\beta^{*})^{T}(D_{\beta^{*},\beta^{*}}-D_{\beta^{*},\alpha^{*}}D^{-1}_{\alpha^{*},\alpha^{*}}D_{\alpha^{*},\beta^{*}})(\hat{\beta}-\beta^{*})+O_{p}(n^{-3/2})$ (A.7)
+[…]
 However, $\log P_{n}\{W(x,p_{M(\hat{\alpha},\hat{\beta})})\}-\log P_{n}\{W(x,p_{M(\hat{\alpha}_{0},\beta^{*})})\}=\log GIC(k)-\log GIC(k_{0}).$ Recall $\sqrt{n}(\hat{\theta}-\theta^{*})$ is asymptotically normally distributed with mean zero and covariance matrix equal to $D^{-1}(\theta^{*})\Lambda(\theta^{*})D^{-T}(\theta^{*})$.
@@
 Hence, $n\times[\log GIC(k)-\log GIC(k_{0})]$ converges in distribution to $Z^{T}A(\theta^{*})Z$.
-A.2 Proof of Theorem 1
+∎
+[…]
 Proof.
@@
 Then, $\mathcal{M}_{k}<\mathcal{M}_{k_{0}}$ for all $1\leq k<k_{0}$, whereas $\mathcal{M}_{k}=\mathcal{M}_{k_{0}}$ otherwise.
-Note that from Proposition 3 in Cheng and Tong (2024), $\mathcal{M}_{k_{0}}=H_{G}(p^{*})>0$ where $p^{*}$ is the true population pdf.
-Therefore, if for any $k$, $C(n,k)\to 1$ as $n\to\infty$, then the maximum T-GIC model selection criterion will not select any $k<k_{0}$, in probability.
+Note that from Proposition 3 in Cheng & Tong (2024), $\mathcal{M}_{k_{0}}=H_{G}(p^{*})>0$ where $p^{*}$ is the true population pdf.
+Therefore, if for any $k$, $C(n,k)\to 1$ as $n\to\infty$, then the maximum MIC model selection criterion will not select any $k<k_{0}$, in probability.
 Henceforth, consider the case that $k\geq k_{0}$.
@@
 Let $k>k_{0}$ be fixed.
-Consider the increment $D=\log\{T\text{-}GIC(k)\}-\log\{T\text{-}GIC(k_{0})\}=\log C(n,k)-\log C(n,k_{0})+\log GIC(k)-\log GIC(k_{0})$.
+Consider the increment $D=\log\{MIC(k)\}-\log\{MIC(k_{0})\}=\log C(n,k)-\log C(n,k_{0})+\log GIC(k)-\log GIC(k_{0})$.
 By Proposition 1, $n\times\{\log GIC(k)-\log GIC(k_{0})\}$ converges weakly to some non-negative distribution.
 Consequently, $D=\log C(n,k)-\log C(n,k_{0})+O_{p}(1/n)$ so that if $n\times\log\{C(n,k)/C(n,k_{0})\}\to$-$\infty$ as $n\to\infty$, $D$ is negative in probability for $k>k_{0}$.
-This completes the proof of the consistency of the proposed T-GIC model selection criterion.
+This completes the proof of the consistency of the proposed MIC model selection criterion.
+∎